### v1

# Finite–slope universal Rankin–Selberg $p$-adic $L$-function

Abstract
In this article we construct a finite–slope universal Rankin–Selberg $p$-adic $L$-function attached to a half–ordinary universal deformation family of $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ Galois representations. More precisely, starting from residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ of tame level $1$ satisfying Hypothesis 3.1 of Loeffler [13], we consider the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of [13], where the first factor varies in the ordinary Hida deformation and the second factor in the unrestricted universal deformation space. Fixing a classical point $x_{0}$ on a suitable parabolic eigenvariety for which the first factor is non–ordinary but of small finite slope, we use Liu’s global triangulation theorem, Kedlaya–Pottharst–Xiao’s $(\varphi,\Gamma)$–cohomology, and the Perrin–Riou–Loeffler–Zerbes regulator formalism to attach to the resulting Panchishkin family $(V_{U},V_{U}^{+})$ over a neighbourhood $U$ of $x_{0}$ a “big logarithm”
$\mathcal{L}_{V_{U},V_{U}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)\cong\mathcal{O}(U\times\mathscr{W}),$
interpolating the Bloch–Kato dual exponentials at all classical points in $U$.
Combining this family regulator with the universal Beilinson–Flach Euler system of Loeffler-Zerbes[21], as interpreted in the universal deformation setting in [15], we define a rigid–analytic function
$L_{p}^{\mathrm{fs}}:=\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})\in\mathcal{O}(U\times\mathscr{W}),$
and prove that $L_{p}^{\mathrm{fs}}$ satisfies the expected interpolation formula for all Deligne–critical values of the complex Rankin–Selberg $L$-functions $L(f_{x}\otimes g_{x},s)$ at classical points $(x,\kappa)\in U\times\mathscr{W}$. In particular, under mild crystalline and non–criticality hypotheses at $p$, we verify the finite–slope analogue of Loeffler’s Conjecture 2.8 in the concrete $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ half–ordinary universal deformation setting of [13].
The main new features of this work are: (1) the construction of a Perrin–Riou regulator for a genuinely finite–slope Panchishkin family over a parabolic eigenvariety in the universal deformation context, and (2) the resulting finite–slope universal Rankin–Selberg $p$-adic $L$-function, which simultaneously extends the ordinary universal results of Loeffler and Hao–Loeffler [13, 15] and the finite–slope eigencurve constructions of Loeffler’s earlier three–variable Rankin–Selberg $p$-adic $L$-functions. This provides the first verification of the universal finite–slope conjectural picture in the Rankin–Selberg case and offers a conceptual bridge between ordinary Hida-theoretic constructions and finite–slope Coleman–style $p$-adic $L$-functions.
1 Introduction
1. Global goal and unsolved question
We fix a prime $p\geq 3$ and an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. Let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and $G_{\mathbb{Q},\{p\}}\subset G_{\mathbb{Q}}$ denote the Galois group of the maximal extension of $\mathbb{Q}$ unramified outside $p$.
Let
$\bar{\rho}_{1},\bar{\rho}_{2}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(\mathbb{F}_{p})$
be continuous, odd, absolutely irreducible Galois representations arising from cuspidal newforms $f_{0},g_{0}$ of weights $k_{1},k_{2}\geq 2$ and tame level $N=1$ (unramified outside $p$), so that Hypothesis 3.1 of [13] applies. In particular (see [13, Hyp. 3.1]), we assume that:
  • both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ are modular of level $1$ and weight $\geq 2$;
  • the usual Taylor–Wiles hypotheses hold (absolute irreducibility, oddness, and minimality at auxiliary primes);
  • at $p$ the representation $\bar{\rho}_{1}$ is ordinary and admits a fixed ordinary refinement in the sense of [13, Def. 3.11].
We henceforth keep this tame level $N=1$ fixed throughout.
  • Let $R_{1}$ be the ordinary Hida Hecke algebra attached to $\bar{\rho}_{1}$, and let
    $\mathscr{X}_{1}:=\operatorname{Spf}(R_{1})^{\operatorname{rig}}$
    be the corresponding ordinary eigenvariety, as in [13, §3.2, Prop. 3.14], building on the ordinary families of Hida and the $R=\mathbb{T}$ theorems of Mazur, Böckle and Emerton.
  • Let $R_{2}$ be the universal deformation ring of $\bar{\rho}_{2}$ (for deformations unramified outside $p$) and let $(R_{2},V_{2})$ be the Galois deformation family considered in [13, Def. 3.3, Thm. 3.4] and in [14, 15]. We denote
    $\mathscr{X}_{2}:=\operatorname{Spf}(R_{2})^{\operatorname{rig}}$
    for the associated rigid analytic space.
Following [13, Ex. 3.17], there is a Panchishkin family $(R,V,V^{+})$ for the Rankin–Selberg tensor
$V:=V_{1}^{\mathrm{ord}}\otimes V_{2}$
over
$R:=R_{1}\,\widehat{\otimes}_{\mathbb{Z}_{p}}\,R_{2},$
where $V_{1}^{\mathrm{ord}}$ is the universal ordinary deformation of $\bar{\rho}_{1}$ and $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ is the rank–$1$ ordinary submodule at $p$. More precisely, there is a rank–$r$ locally free RR–submodule
$V^{+}\;\subset\;V\big|_{G_{\mathbb{Q}_{p}}}$
[…]
Loeffler’s Conjecture 2.8 [13, Conj. 2.8] predicts the existence of a rank 0 Euler system (i.e. a $p$-adic $L$-function) attached to $(R,V,V^{+})$; in the ordinary setting this is now known by [15, Thm. 3.5], building on Urban’s construction via nearly overconvergent forms [7] and Hida’s earlier work on Rankin–Selberg pp–adic LL–functions [1, 2].
In [13, §5.5] Loeffler sketches a finite-slope analogue of this conjecture on the Coleman–Mazur eigencurve, and in [13, §6] he formulates related conjectures over big parabolic eigenvarieties (see also [9, 12]). However, beyond very special cases for $\operatorname{GL}_{2}$–families and the rank-one case of [11], no general finite-slope result is currently known.
Unsolved question (finite-slope universal Rankin–Selberg). Let $\mathscr{E}_{1}$ be the Coleman–Mazur eigencurve (or more generally the big parabolic eigenvariety of Barrera Salazar–Williams) attached to $\bar{\rho}_{1}$ [23, 9], and consider the product eigenvariety
$\mathscr{E}\;:=\;\mathscr{E}_{1}\times\mathscr{X}_{2}$
together with the natural weight map to the pp–adic weight space $\mathscr{W}$ and the global Galois family $(V,V^{+})$ coming from [13, Ex. 3.17].
Given a classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ of cuspidal eigenforms with $f_{0}$ of non-ordinary finite slope at $p$, does there exist a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on a neighbourhood $U$ of $x_{0}$ whose specialisations interpolate all Deligne-critical Rankin–Selberg values $L(f_{x}\otimes g_{x},s_{x})$ for classical points $x\in U$?
This is a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ instance of the finite-slope variants of [13, Conj. 2.8] over parabolic eigenvarieties.
To treat this question we need a precise description of the eigenvariety $\mathscr{E}$ near $x_{0}$. We recall the relevant geometric input from [9, Han17], in a form adapted to our setting.
Lemma 1 (Good neighbourhood on the parabolic eigenvariety).
Let $\mathscr{E}$ be the parabolic eigenvariety of [9] with weight space $\mathscr{W}$ and weight map $w:\mathscr{E}\to\mathscr{W}$. Let $x_{0}\in\mathscr{E}$ be a point corresponding to a cuspidal automorphic representation of regular (cohomological) weight, and fix a pp–refinement which is QQ–non-critical in the sense of [9, Def. 5.13]. Assume that the derived group $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series.
Then there exists an affinoid neighbourhood $U\subset\mathscr{E}$ of $x_{0}$ such that:
  1. (i)
    $\mathscr{E}$ is smooth over $\mathbb{Q}_{p}$ at every point of $U$;
  2. (ii)
    the restriction $w|_{U}:U\to w(U)\subset\mathscr{W}$ is finite étale; and
  3. (iii)
    classical cuspidal points are Zariski dense in $U$.
Proof.
By [9, Def. 5.11] and [9, Rem. 1.4], the overconvergent defect $\ell_{Q}(x_{0})$ vanishes at a QQ–non-critical point; in particular $x_{0}$ lies in the interior locus in the sense of [9, Def. 5.11]. Hence [9, Prop. 5.12] implies that every irreducible component $V$ of $\mathscr{E}$ containing $x_{0}$ has dimension at least $\dim\mathscr{W}$.
On the other hand, $\mathscr{E}$ is constructed as a closed subspace of the universal eigenvariety of Hansen [Han17]. By [Han17, Thm. 4.5.1(i)], applied to the corresponding classical, non-critical interior point of the universal eigenvariety, every irreducible component of the latter has dimension equal to the weight space. Since $\mathscr{E}$ is a closed subspace of this universal eigenvariety, every irreducible component $V$ of $\mathscr{E}$ through $x_{0}$ has dimension at most $\dim\mathscr{W}$. Thus $\dim V=\dim\mathscr{W}$ for each component $V$ passing through $x_{0}$.
Under the same hypotheses, [9, Def. 5.13 and Cor. 5.16] show that on each such component $V$ the classical cuspidal points are Zariski dense.
The weight map $w:\mathscr{E}\to\mathscr{W}$ is finite by [9, Thm. 5.4]. Since the source and the target have the same local dimension at $x_{0}$, the morphism $w$ is unramified at $x_{0}$. For finite morphisms between equidimensional rigid analytic spaces, finiteness and unramifiedness imply that $w$ is étale at $x_{0}$ and that $\mathscr{E}$ is smooth at $x_{0}$. The smooth locus of $\mathscr{E}$ and the étale locus of $w$ are open subsets of $\mathscr{E}$, so we may choose an affinoid neighbourhood $U$ of $x_{0}$ contained in their intersection. By the density of classical points on each component through $x_{0}$, the intersection of $U$ with the classical cuspidal locus is Zariski dense in $U$. This proves the lemma. ∎
Remark 1 (On the discrete-series hypothesis in Lemma 1).
In our Rankin–Selberg situation we work with
$G\;=\;\mathrm{GL}_{2}\times\mathrm{GL}_{2},$
so that
$G_{\mathrm{der}}(\mathbb{R})\;=\;\mathrm{SL}_{2}(\mathbb{R})\times\mathrm{SL}_{2}(\mathbb{R}),$
which is well known to admit discrete series. Thus the hypothesis that $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series in Lemma 1 is automatically satisfied in our setting.
The only role of this hypothesis is to allow us to invoke the eigenvariety results of [9, Han17], in particular the dimension and density statements of [9, Prop. 5.12, Cor. 5.16] and [Han17, Thm. 4.5.1(i)], in order to construct an affinoid neighbourhood $U$ of $x_{0}$ which is smooth, finite étale over weight space, and has Zariski-dense classical cuspidal locus. All later arguments only use the existence of such a neighbourhood $U$ and do not appeal directly to the discrete-series condition.
2. Target main theorem (precise special case)
Fix a finite-slope classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ with:
  • $f_{0}$ a $p$-stabilised newform of weight $k_{1}\geq 2$ and level prime to $p$, with $U_{p}$–eigenvalue $\alpha_{f_{0}}$ of slope
    $0<v_{p}(\alpha_{f_{0}})<k_{1}-1,$
    so that $f_{0}$ is non-ordinary of small slope at $p$;
  • $g_{0}$ the specialisation at a classical arithmetic point of the universal deformation family $(R_{2},V_{2})$, of weight $k_{2}\geq 2$, such that $g_{0}$ is ordinary at $p$ (or more generally satisfies a suitable Panchishkin condition as in [13, Def. 2.1, Def. 2.3] and [15, §2]).
Let $U\subset\mathscr{E}$ be the affinoid neighbourhood of $x_{0}$ given by Lemma 1. Thus $\mathscr{E}$ is smooth at every point of $U$, the weight map $w$ is finite étale on $U$, and classical cuspidal points are Zariski dense in $U$.
After shrinking $U$ if necessary, we may and do assume that:
  • the eigenvariety is smooth on $U$ and the weight map $w\colon U\to w(U)\subset\mathscr{W}$ is finite étale (Lemma 1);
  • the slopes of the $U_{p}$-eigenvalues of $f_{x}$ remain $k_{1}(x)-1$ for all $x\in U$, so that overconvergent classicality holds for $f_{x}$ by Coleman [22, Thm. 6.1] (in the form of his Theorem 8.3: weight $k+2$ and slope <k+1<k+1 implies classicality);
  • the Panchishkin submodule $V^{+}$ extends as a rank–$r$ subbundle of the relative $(\varphi,\Gamma)$–module attached to $V$ over $U$, as in [13, Def. 2.1, Def. 2.3, Ex. 3.17] and Liu’s global triangulation theorem [20, Thm. 1.8].
Conjecture 2 (Finite-slope universal Rankin–Selberg $p$-adic $L$-function).
There exists a rigid-analytic function
$L_{p}^{\mathrm{fs}}\;\in\;\mathscr{O}(U\times\mathscr{W})$
on $U$ times the $p$-adic cyclotomic weight space $\mathscr{W}$ such that for each classical point $(x,\kappa)\in U\times\mathscr{W}$ corresponding to a pair $(f_{x},g_{x})$ and an integer critical value $s=\kappa(x)$ of $L(f_{x}\otimes g_{x},s)$ (in Deligne’s sense), we have an interpolation formula of the form
$L_{p}^{\mathrm{fs}}(x,\kappa)\;=\;\frac{E_{p}(f_{x},g_{x},s)}{\Omega_{p}(f_{x},g_{x},\pm)}\cdot\frac{L^{(p)}(f_{x}\otimes g_{x},s)}{(2\pi i)^{2s}},$
where:
  • $E_{p}(f_{x},g_{x},s)$ is the explicit local Euler factor at $p$ appearing in [15, Def. 3.4];
  • $\Omega_{p}(f_{x},g_{x},\pm)$ is a $p$-adic period depending analytically on $x$ (and on the choice of sign $\pm$) and normalised compatibly with [15];
  • $L^{(p)}(f_{x}\otimes g_{x},s)$ is the complex Rankin–Selberg $L$-function with the Euler factor at $p$ omitted.
Theorem 3 (Target main result of the paper).
Assume:
  1. (H1)
    The residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ satisfy Hypothesis 3.1 of [13] (Taylor–Wiles conditions, local conditions at $p$, and tame level $1$). In particular, both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ arise from cuspidal newforms of level $1$ and weights $\geq 2$, and we fix once and for all this tame level $N=1$ throughout. Moreover, $\bar{\rho}_{1}$ admits an ordinary refinement at $p$ in the sense of [13, Def. 3.11].
  1. (H2)
    the point $x_{0}$ is crystalline at $p$ and satisfies the small-slope and non-criticality hypotheses needed for overconvergent classicality (in the sense of Coleman [22, Thm. 6.1]);
  2. (H3)
    the Panchishkin condition holds for $(V,V^{+})$ on $U$ in the sense of [13, Def. 2.1, Def. 2.3] (in particular, the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of loc. cit. uniformly over $U$).
Then Conjecture 2 holds for the neighbourhood $U$. In particular, there exists a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on $U\times\mathscr{W}$ satisfying the above interpolation formula for all classical points in $U$.
The rest of the paper is devoted to the construction of $L_{p}^{\mathrm{fs}}$ and the proof of Theorem 3. Each step is explicitly referenced to the corresponding statements in [13, 19, 20, 16, 21, 25, 15, 14] so that the argument can be checked line by line.
3. New features of this work, obstacles, and future directions.
We conclude the introduction by explaining how Theorem 3 fits into the existing conjectural picture of universal Rankin–Selberg $p$-adic $L$-functions, why the finite–slope case treated here has remained out of reach so far, and which directions seem most promising for further developments.
3.1. Relation to Loeffler’s conjectural framework and to Hao–Loeffler.
In his work on $p$-adic $L$-functions in universal deformation families, Loeffler formulates a general conjecture (Conjecture 2.8 of [13]) predicting the existence and interpolation properties of rank 0 Euler systems (i.e. $p$-adic $L$-functions) attached to 0-Panchishkin families $(R,V,V^{+})$ over suitable “big parabolic eigenvarieties”. His main examples include the half–ordinary Rankin–Selberg tensor considered in Example 3.17 of loc. cit., which is precisely the global deformation setting adopted in Step 1 of this article.
In the ordinary case, this conjectural picture has now been confirmed for the Rankin–Selberg tensor by the recent work of Hao–Loeffler [15]. Building on Urban’s nearly overconvergent three–variable Rankin–Selberg $p$-adic $L$-function [7] and the Beilinson–Flach Euler system of Loeffler–Zerbes [21], they construct a universal Rankin–Selberg $p$-adic $L$-function for an ordinary Hida family tensored with a universal deformation family, and prove that it interpolates all Deligne–critical Rankin–Selberg values in this setting, together with a functional equation.
However, both [13] and [15] work under a global ordinarity (or nearly–ordinarity) hypothesis at $p$, and the $p$-adic $L$-functions they construct live naturally over ordinary eigenvarieties or ordinary loci in universal deformation spaces. In contrast, in the present paper we fix a classical point $x_{0}=(f_{0},g_{0})$ at which the $f_{0}$–factor is of non–ordinary finite slope and we work in a neighbourhood $U$ of $x_{0}$ in a parabolic eigenvariety in the sense of Barrera Salazar–Williams [9]. Theorem 3 should therefore be viewed as a first verification, in a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ situation, of the finite–slope variants of Loeffler’s Conjecture 2.8 over big parabolic eigenvarieties suggested in [13, §5.5, §6].
3.2. Comparison with existing finite–slope constructions.
Finite–slope $p$-adic $L$-functions attached to families of automorphic forms have been constructed in several other settings, but always in a way that is complementary to the present work.
  • For Rankin–Selberg convolutions of modular forms, Urban’s three–variable $p$-adic $L$-function is constructed on eigenvarieties using nearly overconvergent modular forms and overconvergent cohomology [7], and requires a nearly ordinary hypothesis at $p$. These functions live over eigenvarieties and are not formulated in terms of universal Galois deformation spaces.
  • The geometric methods of Andreatta–Iovita and their collaborators yield triple product $p$-adic $L$-functions for finite slope families of modular forms via the theory of overconvergent sheaves and the spectral halo [11]. Again, the base spaces are eigenvarieties, and the Galois representations are not packaged as a single Panchishkin family $(R,V,V^{+})$ over a universal deformation ring.
  • Barrera Salazar–Dimitrov–Williams construct finite–slope $p$-adic $L$-functions for Shalika families on $\operatorname{GL}_{2n}$ over parabolic eigenvarieties, and relate their existence to the local geometry of these eigenvarieties [12]. Their methods are adapted to the standard representation of $\operatorname{GL}_{2n}$ and do not provide a universal deformation interpretation for Rankin–Selberg tensors of $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$.
By contrast, the present article keeps the universal deformation viewpoint of [13, 15], starting from the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of [13], and then passes to a finite–slope neighbourhood $U$ on the parabolic eigenvariety through the identification $U\simeq U_{R}\subset\operatorname{Spf}(R)_{\mathrm{rig}}$ of Step 1.4. The resulting finite–slope universal Rankin–Selberg $p$-adic $L$-function
$L_{p}^{\mathrm{fs}}\;=\;\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})\in\mathcal{O}(U\times\mathcal{W})$
simultaneously interpolates the critical Rankin–Selberg values for all classical points in $U$ and specialises to the ordinary universal Rankin–Selberg $p$-adic $L$-function at ordinary points. In this way it provides a conceptual bridge between the ordinary universal results of [13, 15] and the finite–slope eigenvariety constructions along the lines of Urban and Andreatta–Iovita [7, 11].
3.3. Why the finite–slope case is difficult.
From the perspective of [13], it is natural to expect that a finite–slope analogue of Conjecture 2.8 should hold whenever one can attach a $(\varphi,\Gamma)$-module to the Panchishkin family $(R,V,V^{+})$ and construct an appropriate Perrin–Riou regulator [18]. However, there are several serious obstacles which have so far prevented a general theory, and which Theorem 3 overcomes in the specific Rankin–Selberg case treated here:
[…]
  • Cohomology and base change. Even after triangulation, one must control the $(\varphi,\Gamma)$- and Iwasawa cohomology of $D$ and $D^{+}$ in families, with good base–change and perfectness properties, in order to set up a Perrin–Riou type regulator over the affinoid algebra $\mathcal{O}(U)$. This is achieved here by systematically using the cohomological machinery of Kedlaya–Pottharst–Xiao [19], which guarantees that the relevant cohomology complexes lie in the perfect derived category and behave well under specialisation.
  • Family Perrin–Riou regulator for a genuine finite–slope Panchishkin family. Existing constructions of big logarithms and regulators in families (following Perrin–Riou, Coleman, Loeffler–Zerbes, Nakamura, Pottharst, etc.; see for instance [18, 21, 16]) either treat a single de Rham representation or work under an ordinarity hypothesis. The map
    $\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\;\longrightarrow\;\mathcal{H}(\Gamma)\,\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$
    constructed in Step 3 is, to the best of our knowledge, the first instance of a Perrin–Riou regulator attached to a genuinely finite–slope Panchishkin family over a higher–dimensional parabolic eigenvariety in the universal deformation setting.
  • Compatibility with Beilinson–Flach Euler systems. Finally, one must show that the universal Beilinson–Flach classes of Loeffler–Zerbes [21], interpreted in the universal deformation setting (following [15]), specialise correctly along $U$ and lie in the critical Iwasawa cohomology group on which $\mathcal{L}_{V_{U},V_{U}^{+}}$ is defined. This compatibility is crucial in order to define $L_{p}^{\mathrm{fs}}$ and deduce its interpolation formula from the explicit reciprocity laws for Beilinson–Flach elements.
These issues explain why, despite the conceptual clarity of the finite–slope variant of Conjecture 2.8, no general finite–slope universal Rankin–Selberg $p$-adic $L$-function had been constructed before.
3.4. Future directions.
The techniques developed in this article suggest several natural directions for further work:
  • Beyond the small–slope, non–critical range. A first goal would be to relax the crystalline small–slope and non–criticality hypotheses at $p$ imposed in Theorem 3. This would require a finer analysis of triangulations and local Galois representations at critical slope, possibly using the spectral–halo techniques of Andreatta–Iovita and the overconvergent cohomology methods of Barrera Salazar–Williams [11, 9].
  • Other groups and higher rank. It should be possible, at least in principle, to adapt the strategy of Steps 1–3 to other instances of Loeffler’s conjectural framework, for example to Shalika families on $\operatorname{GL}_{2n}$ or to $\mathrm{GSp}_{4}$-type Rankin–Selberg convolutions, combining the parabolic eigenvarieties of Barrera Salazar–Williams [9, 12] with the Euler systems and Perrin–Riou regulators available in those settings.
  • Iwasawa–theoretic applications and main conjectures. The finite–slope universal Rankin–Selberg $p$-adic $L$-function $L_{p}^{\mathrm{fs}}$ should control Selmer groups for the specialisations $V_{x}$ in a manner predicted by the Bloch–Kato conjectures and noncommutative Iwasawa theory. One long–term objective is to formulate and prove main conjectures relating characteristic ideals of universal Selmer modules over $R$ to the ideal generated by $L_{p}^{\mathrm{fs}}$.
  • Functional equations and exceptional zeros. The universal functional equation obtained in [15] for the ordinary Rankin–Selberg $p$-adic $L$-function suggests that an analogue should exist for $L_{p}^{\mathrm{fs}}$, relating its values (or derivatives) at critical points to those of other $p$-adic $L$-functions. Understanding such a functional equation in the finite–slope setting could lead to new results on exceptional zeros and on the variation of local epsilon factors in families.
  • Refined arithmetic information. Finally, one may hope to extract finer arithmetic invariants from $L_{p}^{\mathrm{fs}}$, such as $p$-adic heights or regulators on motivic cohomology classes varying in families, in the spirit of the work of Darmon–Rotger and Hsieh on triple product $p$-adic $L$-functions. This would provide further evidence that big parabolic eigenvarieties are indeed the correct parameter spaces for Euler systems and $p$-adic $L$-functions beyond the ordinary setting.
In summary, Theorem 3 provides the first finite–slope verification of Loeffler’s universal conjectural picture in the Rankin–Selberg case, and the methods developed here should be of independent interest for the study of $p$-adic $L$-functions and Euler systems on parabolic eigenvarieties beyond the ordinary locus.
2 Proof and Results
Step 1. Global deformation space and half–ordinary Panchishkin family
We briefly recall the deformation-theoretic input and the construction of the half–ordinary Panchishkin family $(R,V,V^{+})$, following [13, 15]. Recall that $p>2$ is fixed, and that we have two absolutely irreducible residual Galois representations
$\bar{\rho}_{i}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(\mathbb{F}),\qquad i=1,2,$
arising from normalised cuspidal Hecke eigenforms $f_{0}$ and $g_{0}$ of tame level $1$ (or more generally of fixed tame level prime to $p$), with $\bar{\rho}_{1}$ ordinary at $p$ and $\bar{\rho}_{2}$ subject to the usual Taylor–Wiles hypotheses. We write $\mathcal{O}$ for a finite extension of $\mathbb{Z}_{p}$ containing the Hecke eigenvalues of $f_{0}$ and $g_{0}$, and we let $\mathrm{CNL}_{\mathcal{O}}$ denote the category of complete Noetherian local $\mathcal{O}$–algebras with residue field $\mathbb{F}$.
(1.1) Ordinary universal deformation ring for $\bar{\rho}_{1}$.
Let $R^{\mathrm{ord}}(\bar{\rho}_{1})$ denote the universal deformation ring parametrising ordinary deformations of $\bar{\rho}_{1}$ as a $G_{\mathbb{Q},\{p\}}$–representation in the sense of Hida and Mazur: for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation
$\rho_{1,A}:G_{\mathbb{Q},\{p\}}\to\operatorname{GL}_{2}(A)$
is called ordinary if its restriction to $G_{\mathbb{Q}_{p}}$ fits into a short exact sequence
$0\longrightarrow V_{1,A}^{\mathrm{ord},+}\longrightarrow V_{1,A}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V_{1,A}^{\mathrm{ord},-}\longrightarrow 0$
where $V_{1,A}^{\mathrm{ord},+}$ is a rank–$1$ direct summand on which $G_{\mathbb{Q}_{p}}$ acts by an unramified character lifting the pp–adic unit root of the Hecke polynomial of $f_{0}$.
By Mazur’s deformation theory together with Hida theory and the work of Böckle and Emerton (see, for example, [13, §3.2, Prop. 3.14 and Thm. 3.4]), this ordinary deformation problem is representable by a complete Noetherian local $\mathcal{O}$–algebra $R^{\mathrm{ord}}(\bar{\rho}_{1})$, and there exists a universal ordinary deformation
$\rho_{1}^{\mathrm{ord}}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(R^{\mathrm{ord}}(\bar{\rho}_{1}))$
with underlying free rank–$2$ $R^{\mathrm{ord}}(\bar{\rho}_{1})$–module $V_{1}^{\mathrm{ord}}$.
We denote by $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ the rank–$1$ direct summand giving the ordinary filtration at $p$, and by $V_{1}^{\mathrm{ord},-}$ the corresponding quotient. The ring $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is finite flat over the weight space and is canonically isomorphic to the localised ordinary Hecke algebra $T^{\mathrm{ord}}(\bar{\rho}_{1})$ acting on ordinary pp–adic modular forms of tame level $1$; in particular $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is reduced and equidimensional of relative dimension $1$ over $\mathcal{O}$ (see [13, §3.2, Prop. 3.14]).
For brevity we set
$R_{1}:=R^{\mathrm{ord}}(\bar{\rho}_{1}),\qquad V_{1}^{\mathrm{ord}}\text{ and }V_{1}^{\mathrm{ord},+}\text{ as above.}$
(1.2) The unrestricted universal deformation ring and representation for $\bar{\rho}_{2}$.
For the second factor, we do not impose any ordinary or nearly–ordinary local condition at $p$. We let $R(\bar{\rho}_{2})$ be the universal deformation ring parametrising deformations of $\bar{\rho}_{2}$ as a $G_{\mathbb{Q},\{p\}}$–representation (i.e. unramified outside $p$), as in [13, Def. 3.3]. Thus, for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation to $A$ is simply a continuous representation
$G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(A)$
lifting $\bar{\rho}_{2}$ and unramified outside $p$.
Then the main $R=\mathbb{T}$ theorem of Böckle–Emerton in this context (as formulated in [13, Thm. 3.4], corresponding to [13, Thm. 2.4]) shows that:
  • The ring $R(\bar{\rho}_{2})$ is a reduced complete intersection ring, flat over $\mathcal{O}$ of relative dimension $3$.
  • There is a canonical isomorphism
    $R(\bar{\rho}_{2})\;\cong\;T(\bar{\rho}_{2}),$
    where $T(\bar{\rho}_{2})$ is the localisation at the maximal ideal corresponding to $\bar{\rho}_{2}$ of the prime-to-$p$ Hecke algebra acting on the space $S(1,\mathcal{O})$ of cuspidal pp–adic modular forms of tame level $1$.
By definition of the universal deformation, there is a free rank–$2$ $R(\bar{\rho}_{2})$–module $V_{2}$ equipped with a continuous $G_{\mathbb{Q},\{p\}}$–action lifting $\bar{\rho}_{2}$; we regard this as the universal pp–adic Galois representation of type $\bar{\rho}_{2}$.
We henceforth abbreviate
$R_{2}:=R(\bar{\rho}_{2}),\qquad V_{2}\text{ as above.}$
(1.3) The half–ordinary Rankin–Selberg Panchishkin family.
We now combine the two deformation problems. Consider the completed tensor product
$R\;:=\;R_{1}\,\widehat{\otimes}_{\mathcal{O}}\,R_{2}.$
Since $R_{1}$ and $R_{2}$ are flat complete Noetherian local $\mathcal{O}$–algebras of relative dimensions $1$ and $3$ respectively, their completed tensor product $R$ is again a flat complete Noetherian local $\mathcal{O}$–algebra of relative dimension
$\dim R\;=\;\dim R_{1}+\dim R_{2}\;=\;1+3\;=\;4,$
and the residual representation is $\bar{\rho}_{1}\otimes\bar{\rho}_{2}$.
On $R$ we have the rank–$4$ RR–module
$V\;:=\;V_{1}^{\mathrm{ord}}\,\widehat{\otimes}_{\mathcal{O}}\,V_{2}$
with its diagonal $G_{\mathbb{Q}}$–action, and the rank–$2$ RR–submodule
$V^{+}\;:=\;V_{1}^{\mathrm{ord},+}\,\widehat{\otimes}_{\mathcal{O}}\,V_{2}\;\subset\;V\big|_{G_{\mathbb{Q}_{p}}},$
which is stable under $G_{\mathbb{Q}_{p}}$ because both factors are. This gives a $G_{\mathbb{Q}_{p}}$–stable filtration
$0\longrightarrow V^{+}\longrightarrow V\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V^{-}\longrightarrow 0$
with $V^{+}$ of rank $2$ and $V^{-}$ of rank $2$ over $R$.
This is exactly the datum considered in [13, Ex. 3.17], specialised to the case where the first factor is already ordinary, so that no additional twist is needed to make the ordinary filtration at $p$ appear. In the notation of [13, Ex. 3.17], we are taking
$R=R^{\mathrm{ord}}(\bar{\rho}_{1})\,\widehat{\otimes}_{\mathcal{O}}\,R(\bar{\rho}_{2}),\qquad V=V^{\mathrm{ord}}_{1}\,\widehat{\otimes}_{\mathcal{O}}V_{2},\qquad V^{+}=V^{\mathrm{ord},+}_{1}\,\widehat{\otimes}_{\mathcal{O}}V_{2}.$
[…]
Moreover, the interpolation set
[…]
is Zariski-dense in $\operatorname{Spf}(R)^{\operatorname{rig}}$ (again [13, Ex. 3.17]). Thus the triple $(R,V,V^{+})$ provides precisely the global deformation space and Panchishkin data required in the hypotheses of Theorem 3.
(1.4) Relation with our chosen neighbourhood $U$.
We now explain how $(R,V,V^{+})$ is related to the eigenvariety $\mathscr{E}$ introduced in the introduction. By construction of the rigid spaces above, the ordinary deformation space $\operatorname{Spf}(R_{1})^{\operatorname{rig}}$ dominates the ordinary locus of the Coleman–Mazur eigencurve (and more generally the parabolic eigenvariety) attached to $\bar{\rho}_{1}$; this is made precise by the $R=\mathbb{T}$ identification between $R_{1}$ and the localised ordinary Hecke algebra, together with the construction of the ordinary eigenvariety (cf. [13, §3.2, Prop. 3.14], and [15, §2] for the Rankin–Selberg set–up). On the other hand, $\operatorname{Spf}(R_{2})^{\operatorname{rig}}$ is, by definition, the universal deformation space for $\bar{\rho}_{2}$ [13, §3.1].
Hence the product
$X:=\operatorname{Spf}(R)^{\operatorname{rig}}\;=\;\operatorname{Spf}(R_{1})^{\operatorname{rig}}\times\operatorname{Spf}(R_{2})^{\operatorname{rig}}$
maps naturally to our Rankin–Selberg eigenvariety $\mathscr{E}$. The general formalism of parabolic eigenvarieties developed by Barrera Salazar–Williams [9] (see in particular [9, Thm. 5.4]), combined with the Rankin–Selberg construction of [15, §2], shows that this map is locally an isomorphism: there exist affinoid neighbourhoods
$U_{R}\subset X,\qquad U_{E}\subset\mathscr{E}$
of the point corresponding to $(f_{0},g_{0})$ such that $U_{R}\cong U_{E}$ as rigid analytic spaces, and $U_{E}$ is identified with the local piece of the parabolic eigenvariety cut out by the Hecke algebra acting on Rankin–Selberg parahoric overconvergent cohomology (cf. [9, Thm. 5.4], together with [13, Prop. 3.14] and [15, §2]).
Shrinking if necessary, we may (and do) choose the affinoid neighbourhood $U\subset\mathscr{E}$ of our fixed classical point $x_{0}=(f_{0},g_{0})$ in Lemma 1 so that
$U\;\simeq\;U_{R}\;\subset\;X=\operatorname{Spf}(R)^{\operatorname{rig}},$
and we henceforth identify $U$ with this connected affinoid subdomain of $\operatorname{Spf}(R)^{\operatorname{rig}}$.
In particular, every classical point $x\in U$ corresponds to a specialisation $(V_{x},V_{x}^{+})$ of the Panchishkin family $(V,V^{+})$ constructed above, and this specialisation coincides with the Rankin–Selberg Galois representation $V(f_{x})\otimes V(g_{x})$ together with the usual half–ordinary local subspace at $p$ (ordinary on the ff–factor, no local condition on the gg–factor), as in [13, Ex. 3.17] and [15, §2]. Thus, after restricting to $U$, the global deformation space and Panchishkin datum appearing in Theorem 3 are entirely furnished by this half–ordinary Rankin–Selberg example of Loeffler.
Step 2. Global triangulation and perfect cohomology complexes
In this step we extract the local $(\varphi,\Gamma)$–module attached to the Panchishkin family $(R,V,V^{+})$ near our fixed point $x_{0}$, and we record the finiteness and base–change properties of its $(\varphi,\Gamma)$–cohomology, following [20, 19].
2.1. The relative $(\varphi,\Gamma)$–module
Let $X:=\operatorname{Spf}(R)^{\operatorname{rig}}$ be the rigid analytic space attached to $R$, and let $V$ be the 44–dimensional RR–linear representation of $G_{\mathbb{Q}}$ constructed in Step 1. As in [19, §2], one associates to the restriction $V|_{G_{\mathbb{Q}_{p}}}$ a family of $(\varphi,\Gamma)$–modules
$\mathscr{D}\;:=\;D_{\operatorname{rig}}^{\dagger}\bigl(V|_{G_{\mathbb{Q}_{p}}}\bigr),$
which is a locally free module of rank $4$ over the relative Robba ring $\mathcal{R}_{X}$, endowed with commuting semilinear actions of $\varphi$ and $\Gamma:=\operatorname{Gal}(\mathbb{Q}_{p}(\mu_{p^{\infty}})/\mathbb{Q}_{p})$. For any point $x\in X$, we write $\mathscr{D}_{x}$ for the fibre; by construction $\mathscr{D}_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x})$ as a $(\varphi,\Gamma)$–module over the usual Robba ring over $\mathbb{Q}_{p}$.
Remark 2.
The existence, functoriality and base–change properties of $D_{\operatorname{rig}}^{\dagger}$ for families of pp–adic Galois representations are given by [19, Thm. 2.2.17]. Concretely, if $A$ is a $\mathbb{Q}_{p}$–affinoid algebra and $V_{A}$ is a finite projective AA–module with a continuous AA–linear action of $G_{\mathbb{Q}_{p}}$, then there is an associated $(\varphi,\Gamma)$–module $D_{\operatorname{rig}}^{\dagger}(V_{A})$ over the relative Robba ring $\mathcal{R}_{A}(\pi)$ (denoted $\mathbf{D}_{\mathrm{rig}}(V_{A})$ in loc. cit.), and for any morphism $A\to B$ of $\mathbb{Q}_{p}$–affinoid algebras one has a natural base–change isomorphism
$D_{\operatorname{rig}}^{\dagger}(V_{A})\,\widehat{\otimes}_{A}B\;\cong\;D_{\operatorname{rig}}^{\dagger}(V_{A}\otimes_{A}B).$
2.2. Global triangulation à la Liu
By (H1) and (H3), the pair $(V,V^{+})$ defines an rr–Panchishkin family over $R$ in the sense of [13, Def. 2.1, Rem. 2.2(i), Def. 2.3]: at classical points in the interpolation region of loc. cit., the specialisations $V_{x}$ satisfy the rr–Panchishkin condition and $V_{x}^{+}$ is the corresponding Panchishkin subrepresentation. In particular, the Hodge–Tate weight data at $p$ coming from $(V,V^{+})$ endows the family of $(\varphi,\Gamma)$–modules $\mathscr{D}$ with the additional structure making it a weakly refined family in the sense of [20, Def. 1.5] (= Definition 13.1 in loc. cit.). We may therefore apply Liu’s global triangulation theorem.
Theorem 4 (Liu, global triangulation).
Let $X$ be a reduced separated rigid analytic space over $\mathbb{Q}_{p}$, and let $V_{X}$ be a weakly refined family of pp–adic representations of $G_{\mathbb{Q}_{p}}$ over $X$ in the sense of [20, Def. 1.5]. Let $\mathscr{D}_{X}:=D_{\operatorname{rig}}^{\dagger}(V_{X})$ be the associated family of $(\varphi,\Gamma)$–modules. Then there exists a Zariski open and dense subspace $X^{\mathrm{tri}}\subset X$ such that $\mathscr{D}_{X}|_{X^{\mathrm{tri}}}$ admits a global triangulation: that is, a filtration by $(\varphi,\Gamma)$–submodules
$0=\operatorname{Fil}^{0}\mathscr{D}_{X}\subset\operatorname{Fil}^{1}\mathscr{D}_{X}\subset\cdots\subset\operatorname{Fil}^{d}\mathscr{D}_{X}=\mathscr{D}_{X}$
whose graded pieces $\operatorname{gr}^{i}\mathscr{D}_{X}:=\operatorname{Fil}^{i}/\operatorname{Fil}^{i-1}$ are rank–$1$ $(\varphi,\Gamma)$–modules with prescribed parameters. Moreover, the locus of global triangulation contains all regular non–critical points of $X$ [20, Thm. 1.8] (corresponding to Theorem 13.2 of loc. cit.).
Applying Theorem 4 to $V_{X}=V|_{G_{\mathbb{Q}_{p}}}$ and shrinking $X$ around $x_{0}$ if necessary, we may and do assume that $x_{0}\in X^{\mathrm{tri}}$ and that the Panchishkin condition at $x_{0}$ is non–critical in the sense of [20, Def. 5.29]. By hypotheses (H2) and (H3) we assume that $x_{0}$ is crystalline with small slope and that $V_{x_{0}}^{+}$ defines a non–critical Panchishkin subspace (i.e. the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of [13, Def. 2.1]), so that $x_{0}$ is a regular non–critical point in the sense of [20, Def. 5.29]. Thus there exists a Zariski open neighbourhood
$U^{\mathrm{tri}}\subset X^{\mathrm{tri}}$
of $x_{0}$ and a filtration
$0=\operatorname{Fil}^{0}\mathscr{D}\subset\operatorname{Fil}^{1}\mathscr{D}\subset\operatorname{Fil}^{2}\mathscr{D}\subset\operatorname{Fil}^{3}\mathscr{D}\subset\operatorname{Fil}^{4}\mathscr{D}=\mathscr{D}$
of $\mathscr{D}|_{U^{\mathrm{tri}}}$ by $(\varphi,\Gamma)$–submodules, such that each $\operatorname{gr}^{i}\mathscr{D}$ is locally free of rank $1$ over the relative Robba ring over $U^{\mathrm{tri}}$.
Replacing our previously fixed affinoid neighbourhood $U\subset X=\operatorname{Spf}(R)^{\operatorname{rig}}$ (from Lemma 1 and Step 1.4) by a smaller affinoid subdomain if necessary, we may and do assume that
$U\subset U^{\mathrm{tri}}$
and that the above filtration is defined over $U$. We continue to denote this neighbourhood by $U$.
Definition 1.
Over the neighbourhood $U\subset X^{\mathrm{tri}}$, we define
$\mathscr{D}^{+}\;:=\;\operatorname{Fil}^{r}\mathscr{D},$
[…]
Proposition 1 (Compatibility with the Panchishkin submodule).
For every classical point $x\in U$ corresponding to a pair $(f_{x},g_{x})$, the fibre of $\mathscr{D}^{+}$ at $x$ is canonically identified with the Panchishkin submodule of $D_{\operatorname{rig}}^{\dagger}(V_{x})$:
$(\mathscr{D}^{+})_{x}\;\cong\;D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})\subset D_{\operatorname{rig}}^{\dagger}(V_{x}).$
Proof.
By the construction of the rr–Panchishkin family $(R,V,V^{+})$ in [13, §2], the submodule $V^{+}\subset V|_{G_{\mathbb{Q}_{p}}}$ is locally free of rank $r$ over $R$ and, for each classical point $x$ in the interpolation region of loc. cit., the fibre $V_{x}^{+}\subset V_{x}$ is the Panchishkin subrepresentation in the sense of [13, Def. 2.1, Rem. 2.2(i), Def. 2.3]. In particular, $V_{x}^{+}$ is the unique $G_{\mathbb{Q}_{p}}$–stable subspace of $V_{x}$ of dimension $\dim V_{x}^{c=+1}-r$ whose Hodge–Tate weights are strictly larger than those of $V_{x}/V_{x}^{+}$, and the associated $(\varphi,\Gamma)$–module $D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})$ is a saturated submodule of $D_{\operatorname{rig}}^{\dagger}(V_{x})$ of rank $r$.
[…]
$(\operatorname{Fil}^{r}\mathscr{D})_{x}\;=\;D_{\operatorname{rig}}^{\dagger}(V_{x}^{+}).$
Since $\mathscr{D}^{+}:=\operatorname{Fil}^{r}\mathscr{D}$ by definition, this proves the claim. ∎
2.3. Finiteness and base change for $(\varphi,\Gamma)$–cohomology
We now record the finiteness and base–change properties of the $(\varphi,\Gamma)$–cohomology and Iwasawa cohomology of $\mathscr{D}$ and $\mathscr{D}^{+}$ over $U$, following [19]. Let us briefly recall the complexes used by Kedlaya–Pottharst–Xiao.
For a $(\varphi,\Gamma)$–module $\mathsf{M}$ over a relative Robba ring $\mathcal{R}_{A}(\pi)$, we denote by $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ the complex computing the usual $(\varphi,\Gamma)$–cohomology, and by $C^{\bullet}_{\psi}(\mathsf{M})$ the complex computing Iwasawa cohomology, as defined in [19, §4].
Theorem 5 (Kedlaya–Pottharst–Xiao).
Let $A$ be an affinoid $\mathbb{Q}_{p}$–algebra and let $\mathsf{M}$ be a $(\varphi,\Gamma)$–module over the relative Robba ring $\mathcal{R}_{A}(\pi)$. Then:
  1. (i)
    The Iwasawa cohomology complex $C^{\bullet}_{\psi}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}^{\infty}_{A}(\Gamma)\bigr)$, and the $(\varphi,\Gamma)$–cohomology complex $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}(A)$ [19, Thm. 4.4.1, Thm. 4.4.2].
  2. (ii)
    For any morphism of affinoid algebras $A\to B$, the natural maps
    $C^{\bullet}_{\psi}(\mathsf{M})\;\otimes^{\mathbf{L}}_{\mathcal{R}^{\infty}_{A}(\Gamma)}\mathcal{R}^{\infty}_{B}(\Gamma)\longrightarrow C^{\bullet}_{\psi}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr),$
    $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})\;\otimes^{\mathbf{L}}_{A}B\longrightarrow C^{\bullet}_{\varphi,\Gamma}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr)$
    are quasi-isomorphisms [19, Thm. 4.4.3].
We apply Theorem 5 with $A=\mathscr{O}(U)$ and $\mathsf{M}=\mathscr{D}$ or $\mathsf{M}=\mathscr{D}^{+}$. This shows that
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}),\;C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}^{+})\;\in\;D^{-}_{\mathrm{perf}}\bigl(\mathscr{O}(U)\bigr),$
while
$C^{\bullet}_{\psi}(\mathscr{D}),\;C^{\bullet}_{\psi}(\mathscr{D}^{+})\;\in\;D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}^{\infty}_{\mathscr{O}(U)}(\Gamma)\bigr).$
Moreover, for any morphism of affinoid algebras $\mathscr{O}(U)\to B$ (for instance $B=\mathscr{O}(U^{\prime})$ for an affinoid subdomain $U^{\prime}\subset U$, or $B=k(x)$ for the residue field at a point $x\in U$), the base–change isomorphisms of Theorem 5(ii) give quasi-isomorphisms
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathscr{O}(U)}B\;\xrightarrow{\ \sim\ }\;C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}\widehat{\otimes}_{\mathscr{O}(U)}B),$
and similarly for $\mathscr{D}^{+}$, as well as
$C^{\bullet}_{\psi}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{R}^{\infty}_{\mathscr{O}(U)}(\Gamma)}\mathcal{R}^{\infty}_{B}(\Gamma)\;\xrightarrow{\ \sim\ }\;C^{\bullet}_{\psi}\bigl(\mathscr{D}\widehat{\otimes}_{\mathscr{O}(U)}B\bigr),$
and analogously for $\mathscr{D}^{+}$. In particular, the formation of these complexes is compatible with restriction to smaller affinoid neighbourhoods and with specialisation at points of $U$.
Remark 3.
For each $x\in U$, the fibre of $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})$ at $x$ computes the $(\varphi,\Gamma)$–cohomology of $D_{\operatorname{rig}}^{\dagger}(V_{x})$, and similarly for $\mathscr{D}^{+}$. Likewise, the fibres of $C^{\bullet}_{\psi}(\mathscr{D})$ and $C^{\bullet}_{\psi}(\mathscr{D}^{+})$ compute the Iwasawa cohomology of $D_{\operatorname{rig}}^{\dagger}(V_{x})$ and $D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})$, respectively. Thus these complexes give a uniform description, over $U$, of the local Galois cohomology groups and the Bloch–Kato local conditions at $p$ for the specialisations $V_{x}$ and $V_{x}^{+}$ in the sense of [16, §3]. This is the cohomological input needed in Step 3 to construct a big logarithm/regulator map
$\mathcal{L}_{V,V^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\;\longrightarrow\;\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Z}_{p}}\mathscr{O}(U),$
and to relate its specialisations at classical points to the Bloch–Kato exponentials appearing in the Bloch–Kato Selmer formalism of [16, §3].
Step 3: Construction of the big logarithm and of the finite–slope Rankin–Selberg regulator
In this step we fix the affinoid neighbourhood $U$ and the Panchishkin family
$(V_{U},V_{U}^{+})$
from Steps 1 and 2. Thus $V_{U}$ is a rank–$4$ finite projective $\mathcal{O}(U)$-module equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$, and for each point $x\in U$ the specialisation $V_{x}$ is de Rham at $p$ and equipped with a Panchishkin subspace $V_{x}^{+}\subset V_{x}$ of fixed rank $r$.
We first introduce Iwasawa–theoretic notation and recall the local Perrin–Riou big logarithm for a single Panchishkin representation. We then explain how to extend it to our family $(V_{U},V_{U}^{+})$ using the relative $(\varphi,\Gamma)$–module technology of [19] and the Perrin–Riou formalism in the guise of [18, 24, 25, 26, 15]. Finally we apply this family big logarithm to the Beilinson–Flach Euler system of [21, 25, 15] to define the finite–slope universal Rankin–Selberg pp–adic LL–function $L_{p}^{\mathrm{fs}}$ and prove its interpolation property.
(3.1) Iwasawa cohomology and the distribution algebra.
Let
$\Gamma:=\operatorname{Gal}\bigl(\mathbb{Q}_{p}(\mu_{p^{\infty}})/\mathbb{Q}_{p}\bigr)\;\cong\;\operatorname{Gal}\bigl(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}\bigr),\qquad\Lambda(\Gamma):=\mathbb{Z}_{p}[[\Gamma]].$
Let
$\mathcal{H}(\Gamma):=\mathcal{H}_{\mathbb{Q}_{p}}(\Gamma)$
denote the algebra of $\mathbb{Q}_{p}$–valued locally analytic distributions on $\Gamma$, i.e. the completion of $\Lambda_{\mathbb{Q}_{p}}(\Gamma)$ in its natural Fréchet topology (cf. [29, §2.2] and [30, §2.1]). Equivalently, $\mathcal{H}(\Gamma)$ is the strong continuous dual of the space of locally analytic $\mathbb{Q}_{p}$–valued functions on $\Gamma$.
For any pp–adic representation $W$ of $G_{\mathbb{Q}_{p}}$ on a finite-dimensional $\mathbb{Q}_{p}$–vector space, its (cyclotomic) Iwasawa cohomology is defined by
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W):=H^{i}\bigl(\mathbb{Q}_{p},W\otimes_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
where $\Lambda(\Gamma)^{\vee}:=\operatorname{Hom}_{\mathrm{cts}}(\Lambda(\Gamma),\mathbb{Q}_{p})$ with the natural $G_{\mathbb{Q}_{p}}$–action; see [18, App. A.2–A.4] and Greenberg [28] for this definition and its basic properties. By construction, $\Lambda(\Gamma)$ acts on $\Lambda(\Gamma)^{\vee}$ via the right regular representation, and hence $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ carries a natural structure of $\Lambda(\Gamma)$-module for every $i$.
Lemma 6 (Finiteness of local Iwasawa cohomology).
If $W$ is a finite-dimensional $\mathbb{Q}_{p}$–representation of $G_{\mathbb{Q}_{p}}$, then $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ is a finitely generated $\Lambda(\Gamma)$–module for all $i\geq 0$. In particular, $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ is finitely generated over $\Lambda(\Gamma)$.
Proof.
Choose a $G_{\mathbb{Q}_{p}}$–stable $\mathbb{Z}_{p}$–lattice $T\subset W$, and write $K_{\infty}=\mathbb{Q}_{p}(\mu_{p^{\infty}})$. Following [18, App. A.2], let
[…]
with respect to the corestriction maps; then $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is naturally a $\Lambda(\Gamma)$–module, and there is a canonical identification
$Z^{i}_{\infty}(\mathbb{Q}_{p},T)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},T):=H^{i}\bigl(\mathbb{Q}_{p},T\otimes_{\mathbb{Z}_{p}}\Lambda(\Gamma)^{\vee}\bigr)$
for all $i$ (see [18, App. A.2–A.3]). Moreover, Proposition A.2.3 of loc. cit. shows that $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is a finitely generated $\Lambda(\Gamma)$–module for $i=0,1,2$, and $Z^{i}_{\infty}(\mathbb{Q}_{p},T)=0$ for $i\geq 3$. Since $W=T\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$, we have
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\;\cong\;Z^{i}_{\infty}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$
as $\Lambda(\Gamma)$–modules, and the finite generation of $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ implies the finite generation of $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ for all $i$. ∎
In order to apply Perrin–Riou’s regulator, we pass from $\Lambda(\Gamma)$–coefficients to the algebra $\mathcal{H}(\Gamma)$ of locally analytic distributions by scalar extension along the canonical map $\Lambda(\Gamma)\to\mathcal{H}(\Gamma)$ and set
$H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)_{\mathcal{H}}\;:=\;H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)\,\widehat{\otimes}_{\Lambda(\Gamma)}\,\mathcal{H}(\Gamma).$
When no confusion can arise, we continue to denote this $\mathcal{H}(\Gamma)$-module simply by $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$.
If $A$ is a $\mathbb{Q}_{p}$–affinoid algebra and $W_{A}$ is a finite projective AA–module with a continuous AA–linear $G_{\mathbb{Q}_{p}}$–action, we define the family Iwasawa cohomology by
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A}):=H^{i}\bigl(\mathbb{Q}_{p},W_{A}\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0);$
this is a priori an $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$–module. Let $\mathcal{R}_{A}^{\infty}(\Gamma)$ denote the Fréchet–Stein Iwasawa algebra of $\Gamma$ over $A$ considered in [19, §4.2]; there is a natural finite flat homomorphism
$A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)\longrightarrow\mathcal{R}_{A}^{\infty}(\Gamma).$
The relative $(\varphi,\Gamma)$–module formalism of Kedlaya–Pottharst–Xiao [19] gives the following more precise statement.
Proposition 2 (Relative local Iwasawa cohomology via $(\varphi,\Gamma)$–modules).
Let $A$ be a reduced $\mathbb{Q}_{p}$–affinoid algebra and $W_{A}$ a finite projective AA–module with a continuous AA–linear $G_{\mathbb{Q}_{p}}$–action. Let $\mathscr{D}_{A}:=D_{\operatorname{rig}}^{\dagger}(W_{A})$ be the associated family of $(\varphi,\Gamma)$–modules over the relative Robba ring $\mathcal{R}_{A}$. Then:
  1. (a)
    there exist functorial complexes
    $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})\in D^{-}_{\mathrm{perf}}(A),\qquad C^{\bullet}_{\psi}(\mathscr{D}_{A})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{A}^{\infty}(\Gamma)\bigr)$
    such that their cohomology groups compute the Galois and Iwasawa cohomology of $W_{A}$ in the sense that
    $H^{i}\bigl(C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})\bigr)\;\cong\;H^{i}(\mathbb{Q}_{p},W_{A}),$
    and
    $H^{i}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})\,\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{R}_{A}^{\infty}(\Gamma)$
    for all $i\geq 0$;
  2. (b)
    the formation of $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, together with the above isomorphisms, commutes with flat base change in $A$;
  3. (c)
    in particular, $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})$ is a finite projective $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$–module, and for every morphism of affinoids $A\to B$ the canonical map
    $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})\,\widehat{\otimes}_{A\widehat{\otimes}\Lambda(\Gamma)}B\widehat{\otimes}\Lambda(\Gamma)\xrightarrow{\ \sim\ }H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{B})$
    is an isomorphism.
Proof.
The existence, perfectness, and base-change properties of the complexes $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, as well as the identifications of their cohomology with Galois and Iwasawa cohomology after extension of scalars to $\mathcal{R}_{A}^{\infty}(\Gamma)$, are proved in [19, Rem. 4.3.3, Prop. 4.3.6, Cor. 4.3.7, Prop. 4.3.8, Thm. 4.4.3]; see also Theorem 5 in Step 2. The finite projectivity of $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})$ over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$ and its base-change property then follow from the fact that $\mathcal{R}_{A}^{\infty}(\Gamma)$ is a Fréchet–Stein algebra, finite flat over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$, and that $H^{1}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)$ is a coadmissible (hence finite projective) $\mathcal{R}_{A}^{\infty}(\Gamma)$–module; see [19, Lem. 4.3.4]. ∎
We apply this proposition with $A=\mathcal{O}(U)$ and $W_{A}=V_{U}^{\ast}(1)$, and write
$H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))$
for the resulting family of local Iwasawa cohomology groups, viewed as a finite projective module over $\mathcal{O}(U)\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$.
(3.2) Classical Perrin–Riou big logarithm.
We recall the local Perrin–Riou regulator for a single de Rham Panchishkin representation.
Theorem 7 (Perrin–Riou big logarithm for a de Rham Panchishkin representation).
Let $V$ be a finite-dimensional $\mathbb{Q}_{p}$–vector space with a continuous de Rham $G_{\mathbb{Q}_{p}}$–action, and let $V^{+}\subset V$ be a Panchishkin subspace in the usual sense (cf. Panchishkin [3] or [16]). Then there exists a canonical $\Lambda(\Gamma)$–linear map
$\mathcal{L}_{V,V^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\otimes_{\mathbb{Q}_{p}}D_{\mathrm{cris}}(V)$
with the following interpolation property: for every integer twist $V(j)$ in the Bloch–Kato range determined by $V^{+}$ and every continuous character $\chi\colon\Gamma\to\overline{\mathbb{Q}}_{p}^{\times}$ of finite order, the specialisation of $\mathcal{L}_{V,V^{+}}$ at $\chi$ is, up to an explicit non-zero scalar depending only on $V$ and $\chi$, the Bloch–Kato dual exponential or logarithm map for $V(j)$.
More precisely, for such $\chi$ there is a commutative diagram
[…]
where $\log^{\ast}_{V(j)}$ is the Bloch–Kato dual exponential or logarithm (depending on the Hodge–Tate weights), and $\operatorname{Fil}^{0}$ is the Hodge filtration on $D_{\mathrm{dR}}(V(j))$.
References for Theorem 7.
The construction and interpolation property are proved in Perrin–Riou’s monograph [18, Ch. 3]. An explicit description via $(\varphi,\Gamma)$–modules, and the comparison with Bloch–Kato exponentials, is given in Berger [24, Thm. II.6]; see also the survey discussion in [16]. ∎
Here $D_{\mathrm{cris}}(V)$ and $D_{\mathrm{dR}}(V)$ denote Fontaine’s filtered $\varphi$–module and de Rham module, respectively.
(3.3) Extension to the family $(V_{U},V_{U}^{+})$.
We now extend Theorem 7 from a single de Rham Panchishkin representation to the family $(V_{U},V_{U}^{+})$ over the affinoid $U\subset X$ constructed in Steps 1 and 2.
Let $A:=\mathcal{O}(U)$ and regard $V_{U}$ as a finite projective AA–module with a continuous AA–linear action of $G_{\mathbb{Q}_{p}}$. As in Step 2, we write
$\mathscr{D}_{U}:=D_{\operatorname{rig}}^{\dagger}(V_{U})$
for the associated family of $(\varphi,\Gamma)$–modules over the relative Robba ring $\mathcal{R}_{A}$, and
$\mathscr{D}_{U}^{+}:=D_{\operatorname{rig}}^{\dagger}(V_{U}^{+})\;\subset\;\mathscr{D}_{U}$
for the saturated submodule corresponding to the Panchishkin subbundle $V_{U}^{+}\subset V_{U}$ (cf. Proposition 1). We set $\mathscr{D}_{U}^{-}:=\mathscr{D}_{U}/\mathscr{D}_{U}^{+}$.
By Theorem 5, there exist functorial complexes
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{U})\in D^{-}_{\mathrm{perf}}(A),\qquad C^{\bullet}_{\psi}(\mathscr{D}_{U})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{A}^{\infty}(\Gamma)\bigr),$
whose cohomology computes the Galois and (cyclotomic) Iwasawa cohomology of $V_{U}$; moreover their formation commutes with flat base change in $A$ [19, Thm. 4.4.1, Thm. 4.4.3].
3.3.1. Crystalline periods for the quotient $V_{U}^{-}$.
The Panchishkin condition (H3) and the global triangulation on $U$ (see Theorem 4 and Proposition 1) imply that for each $x\in U$ the quotient $V_{x}^{-}:=V_{x}/V_{x}^{+}$ is de Rham with all Hodge–Tate weights <0<0, and the $\varphi$–eigenspace
$D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$
for the refined eigenvalue $\alpha_{x}$ is 11–dimensional (cf. [20, Def. 5.29, Rem. 5.30, Prop. 5.31–5.33, Thm. 1.8]). Here $\alpha_{x}$ is the refined $\varphi$–eigenvalue attached to $x$ by the weakly refined family structure.
The next lemma packages the gluing of these eigenspaces into a line bundle of crystalline periods, following Hansen’s construction on the eigencurve.
Lemma 8 (Crystalline period line bundle).
Let $\mathscr{D}_{U}$ and $\mathscr{D}_{U}^{+}$ be as above, and set $\mathscr{D}_{U}^{-}:=\mathscr{D}_{U}/\mathscr{D}_{U}^{+}$. Let $\alpha\in A^{\times}$ be the analytic function giving the refined $\varphi$–eigenvalue on $V_{U}^{-}:=V_{U}/V_{U}^{+}$. Then:
  1. (a)
    the AA–module
    $\mathscr{D}_{\mathrm{cris},U}^{-}:=\bigl(\mathscr{D}_{U}^{-}[1/t]\bigr)^{\Gamma=1,\ \varphi=\alpha}$
    is locally free of rank $1$ on $U$;
  2. (b)
    for each rigid point $x\in U$ we have a canonical identification
    $(\mathscr{D}_{\mathrm{cris},U}^{-})_{x}\;\cong\;D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}.$
Proof.
We divide the argument into three steps.
Step 1: pointwise description. For each rigid point $x\in U$ we have
$\mathscr{D}_{U,x}:=\mathscr{D}_{U}\,\widehat{\otimes}_{A,\kappa(x)}\kappa(x)\;\cong\;D_{\operatorname{rig}}^{\dagger}(V_{x}),$
and similarly $\mathscr{D}_{U,x}^{+}\cong D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})$, by the compatibility of $D_{\operatorname{rig}}^{\dagger}$ with base change [19, Thm. 2.2.17]. Thus
$\mathscr{D}_{U,x}^{-}:=\mathscr{D}_{U}^{-}\,\widehat{\otimes}_{A,\kappa(x)}\kappa(x)\;\cong\;D_{\operatorname{rig}}^{\dagger}(V_{x}^{-}).$
Since $V_{x}^{-}$ is de Rham, Berger’s comparison theorem for $(\varphi,\Gamma)$–modules gives a canonical isomorphism
$D_{\mathrm{cris}}(V_{x}^{-})\;\cong\;\bigl(D_{\operatorname{rig}}^{\dagger}(V_{x}^{-})[1/t]\bigr)^{\Gamma=1},$
compatible with the Frobenius actions (see, for instance, [24, §I.4]). Composing with the identification above we obtain
$D_{\mathrm{cris}}(V_{x}^{-})\;\cong\;\bigl(\mathscr{D}_{U}^{-}[1/t]\bigr)^{\Gamma=1}\otimes_{A,\kappa(x)}\kappa(x).$
By the global triangulation theorem of Liu [20, Thm. 1.8], the quotient $\mathscr{D}_{U}^{-}$ is a trianguline family, and the refined eigenvalue $\alpha_{x}$ is a simple eigenvalue of $\varphi$ on $D_{\mathrm{cris}}(V_{x}^{-})$. Hence the eigenspace $D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$ is 11–dimensional, and it is identified via the preceding isomorphism with
$\bigl(\mathscr{D}_{U}^{-}[1/t]\bigr)^{\Gamma=1,\ \varphi=\alpha}\otimes_{A,\kappa(x)}\kappa(x).$
This will give (b) once we have shown that $\mathscr{D}_{\mathrm{cris},U}^{-}$ is locally free of rank $1$.
Step 2: finiteness and base change. Consider the twist
$\mathscr{D}_{U}^{-,(\alpha^{-1})}:=\mathscr{D}_{U}^{-}\otimes_{A}\mathcal{R}_{A}(\delta),$
where $\delta$ is the rank–$1$ character of $\Gamma\times\{\varphi\}$ sending $\gamma\in\Gamma$ to $1$ and $\varphi$ to $\alpha^{-1}$, and $\mathcal{R}_{A}(\delta)$ is the associated rank–$1$ $(\varphi,\Gamma)$–module over $\mathcal{R}_{A}$. By construction we have
$\bigl(\mathscr{D}_{U}^{-}[1/t]\bigr)^{\Gamma=1,\ \varphi=\alpha}\;=\;\bigl(\mathscr{D}_{U}^{-,(\alpha^{-1})}[1/t]\bigr)^{\Gamma=1,\ \varphi=1}\;=\;H^{0}_{\varphi,\Gamma}\bigl(\mathscr{D}_{U}^{-,(\alpha^{-1})}[1/t]\bigr),$
where $H^{0}_{\varphi,\Gamma}$ denotes the $(\varphi,\Gamma)$–invariants.
For any $(\varphi,\Gamma)$–module $M$ over $\mathcal{R}_{A}$, the cohomology groups $H^{i}_{\varphi,\Gamma}(M)$ are computed by the complex $C^{\bullet}_{\varphi,\Gamma}(M)$ and form finite AA–modules which are compatible with base change in $A$ [19, Thm. 4.4.1, Thm. 4.4.3]. In particular,
$H^{0}_{\varphi,\Gamma}\bigl(\mathscr{D}_{U}^{-,(\alpha^{-1})}[1/t]\bigr)$
is a finite AA–module whose fibre at $x$ is canonically identified with $H^{0}_{\varphi,\Gamma}(\mathscr{D}_{U,x}^{-,(\alpha^{-1})}[1/t])$. For a single representation, the comparison between Galois cohomology and $(\varphi,\Gamma)$–cohomology shows that
$H^{0}_{\varphi,\Gamma}\bigl(D_{\operatorname{rig}}^{\dagger}(V_{x}^{-,(\alpha^{-1})})[1/t]\bigr)\;\cong\;\bigl(D_{\operatorname{rig}}^{\dagger}(V_{x}^{-})[1/t]\bigr)^{\Gamma=1,\ \varphi=\alpha_{x}},$
again by Berger’s results (see, e.g., [24, §I.4]). Thus, by the discussion in Step 1, for each $x$ the fibre of $H^{0}_{\varphi,\Gamma}(\mathscr{D}_{U}^{-,(\alpha^{-1})}[1/t])$ at $x$ is canonically identified with $D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$ and is 11–dimensional over $\kappa(x)$.
Since $A$ is reduced, a finite AA–module whose fibres at all rigid points have dimension $1$ is locally free of rank $1$. Hence
$\mathscr{D}_{\mathrm{cris},U}^{-}\;=\;\bigl(\mathscr{D}_{U}^{-}[1/t]\bigr)^{\Gamma=1,\ \varphi=\alpha}\;=\;H^{0}_{\varphi,\Gamma}\bigl(\mathscr{D}_{U}^{-,(\alpha^{-1})}[1/t]\bigr)$
is a line bundle on $U$, and its fibre at $x$ is identified with $D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$. This proves (a) and (b).
Step 3: relation with Hansen’s construction. The preceding argument is formally identical to Hansen’s construction of the line bundle of crystalline periods on the eigencurve [26, §1.2, Thm. 1.2.2], applied to the quotient $V_{U}^{-}$. In particular, one may view $\mathscr{D}_{\mathrm{cris},U}^{-}$ as the pullback of Hansen’s period line bundle along the natural map from $U$ to the relevant eigenvariety. This provides an alternative proof of (a)–(b) using only the functoriality of the $(\varphi,\Gamma)$–module construction and the perfectness of the KPX cohomology complexes. ∎
After possibly shrinking $U$, we may and do choose a nowhere-vanishing section
$\eta_{U}\in\Gamma\bigl(U,\mathscr{D}_{\mathrm{cris},U}^{-}\bigr),$
so that $\mathscr{D}_{\mathrm{cris},U}^{-}$ is a free AA–module of rank $1$ with basis $\eta_{U}$.
3.3.2. Perrin–Riou regulator in terms of $(\varphi,\Gamma)$–modules.
The construction of $\mathcal{L}_{V,V^{+}}$ in Theorem 7 admits a reinterpretation purely in terms of $(\varphi,\Gamma)$–modules, and this reinterpretation is functorial in the coefficient ring. More precisely, let $(V,V^{+})$ be a de Rham Panchishkin representation, with associated $(\varphi,\Gamma)$–module $D:=D_{\operatorname{rig}}^{\dagger}(V)$ and period line
$D_{\mathrm{cris}}^{-}:=\bigl((D/D^{+})[1/t]\bigr)^{\Gamma=1,\varphi=\alpha},$
where $\alpha$ is the refined Frobenius eigenvalue on $V^{-}:=V/V^{+}$. Then Berger [24] constructs a $\Lambda(\Gamma)$–linear map
$\widetilde{\mathcal{L}}_{V,V^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\longrightarrow\operatorname{Hom}_{\mathbb{Q}_{p}}\bigl(D_{\mathrm{cris}}^{-},\,\mathcal{H}(\Gamma)\bigr)$
whose evaluation at any non-zero vector in $D_{\mathrm{cris}}^{-}$ recovers the classical Perrin–Riou regulator $\mathcal{L}_{V,V^{+}}$. The construction uses only the $(\varphi,\Gamma)$–module $D$ and the crystalline period line $D_{\mathrm{cris}}^{-}$, and is compatible with base change in the coefficient field.
In particular:
  • for a single de Rham Panchishkin representation $(V,V^{+})$, the regulator $\mathcal{L}_{V,V^{+}}$ of Theorem 7 can be constructed purely in terms of $D_{\operatorname{rig}}^{\dagger}(V)$ and the corresponding period line $D_{\mathrm{cris}}^{-}$, as in Berger’s description [24];
  • in the finite–slope setting, analogous regulators in families have been constructed for Coleman and Hida families of modular forms, and more generally over the eigencurve (see Hansen [26, §1.2 and §4.1]). In each case, the key input is the existence of a line bundle of crystalline periods and the functoriality of the $(\varphi,\Gamma)$–module construction.
In our Rankin–Selberg situation over the universal deformation family, the same $(\varphi,\Gamma)$–module formalism, together with the base–change results of [19], yields an AA–linear, $\Lambda(\Gamma)$–linear “vector–valued” regulator for the family $(V_{U},V_{U}^{+})$; this construction is implemented concretely in [15, §3].
Lemma 9 (Vector–valued family regulator).
With notation as above, there exists an AA–linear, $\Lambda(\Gamma)$–linear map
$\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\operatorname{Hom}_{A}\!\bigl(\mathscr{D}_{\mathrm{cris},U}^{-},\,\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}A\bigr),$
characterised by the following properties:
  1. (a)
    for each rigid point $x\in U$, base change along $A\to\kappa(x)$ identifies the fibre of $\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}$ at $x$ with the vector–valued Perrin–Riou regulator $\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}$ attached to $(V_{x},V_{x}^{+})$;
  2. (b)
    for each $x\in U$, evaluating $\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}$ at any non-zero vector in $D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$ recovers the scalar–valued map $\mathcal{L}_{V_{x},V_{x}^{+}}$ of Theorem 7.
Proof.
We sketch the construction and then explain why it is compatible with specialisation.
Construction. By Theorem 5, the Iwasawa cohomology of the family $V_{U}^{\ast}(1)$ can be computed purely in terms of the associated $(\varphi,\Gamma)$–module
$\mathscr{D}_{U}^{\ast}(1):=D_{\operatorname{rig}}^{\dagger}\bigl(V_{U}^{\ast}(1)\bigr),$
via the complex $C^{\bullet}_{\psi}(\mathscr{D}_{U}^{\ast}(1))$ over $\mathcal{R}_{A}^{\infty}(\Gamma)$. In particular,
$H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\,\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{R}_{A}^{\infty}(\Gamma)\;\cong\;H^{1}_{\psi}\bigl(\mathscr{D}_{U}^{\ast}(1)\bigr),$
and this identification is functorial in $A$ and $\Lambda(\Gamma)$.
On the other hand, for a single representation $(V,V^{+})$, Berger expresses Perrin–Riou’s regulator in terms of $(\varphi,\Gamma)$–modules as a morphism of complexes
$C^{\bullet}_{\psi}\bigl(D^{\ast}(1)\bigr)\longrightarrow\operatorname{Hom}\bigl(D_{\mathrm{cris}}^{-},\mathcal{H}(\Gamma)\bigr)[0],$
whose cohomology in degree $1$ is the vector–valued map $\widetilde{\mathcal{L}}_{V,V^{+}}$; see [24, Thm. II.11 and Thm. II.14] and the discussion preceding Theorem 7. This construction depends only on the $(\varphi,\Gamma)$–module $D$ and the line $D_{\mathrm{cris}}^{-}$, and is therefore compatible with base change in the coefficient field.
Specialising Berger’s construction to each fibre $x\in U$, and using the compatibility of $D_{\operatorname{rig}}^{\dagger}$ with base change, we obtain a collection of maps
$\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{x}^{\ast}(1))\longrightarrow\operatorname{Hom}_{\kappa(x)}\bigl(D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}},\,\mathcal{H}(\Gamma)\otimes_{\mathbb{Q}_{p}}\kappa(x)\bigr),$
varying with $x$.
Since $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))$ and $\mathscr{D}_{\mathrm{cris},U}^{-}$ are finite projective over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$ and $A$, respectively (Theorem 5 and Lemma 8), and the construction of $\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}$ is compatible with base change in $x$, there exists a unique AA–linear, $\Lambda(\Gamma)$–linear map
$\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\operatorname{Hom}_{A}\bigl(\mathscr{D}_{\mathrm{cris},U}^{-},\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}A\bigr)$
whose fibre at each $x$ coincides with $\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}$. This is precisely the construction carried out in [15, §3] in the universal Rankin–Selberg setting, where the map is defined at the level of the KPX complexes $C^{\bullet}_{\psi}$ and shown to be compatible with specialisation.
Compatibility with the scalar regulator. For a single representation $(V,V^{+})$, Berger shows that evaluation of $\widetilde{\mathcal{L}}_{V,V^{+}}$ at any nonzero vector in $D_{\mathrm{cris}}^{-}$ recovers the scalar Perrin–Riou regulator $\mathcal{L}_{V,V^{+}}$; this is [24, II.5] and the comparison with Bloch–Kato exponentials used in Theorem 7. Applying this to each fibre $(V_{x},V_{x}^{+})$ and using the identification of $(\mathscr{D}_{\mathrm{cris},U}^{-})_{x}$ with $D_{\mathrm{cris}}(V_{x}^{-})^{\varphi=\alpha_{x}}$ from Lemma 8, we obtain (b). ∎
We now obtain the scalar–valued family regulator by evaluating at the fixed crystalline period $\eta_{U}$.
Proposition 3 (Big logarithm for the family $(V_{U},V_{U}^{+})$).
With notation as above, define
$\mathcal{L}_{V_{U},V_{U}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$
by
$\mathcal{L}_{V_{U},V_{U}^{+}}(z)\;:=\;\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}(z)(\eta_{U})\qquad(z\in H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))).$
Then:
  1. (i)
    $\mathcal{L}_{V_{U},V_{U}^{+}}$ is $\bigl(\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)\bigr)$–linear;
  2. (ii)
    for each rigid point $x\in U$ and each finite-order character $\chi\colon\Gamma\to\overline{\mathbb{Q}}_{p}^{\times}$ in the Panchishkin range, the specialisation of $\mathcal{L}_{V_{U},V_{U}^{+}}$ at $(x,\chi)$ coincides, up to a non-zero scalar depending only on the choice of $\eta_{U}$, with the classical Perrin–Riou regulator of Theorem 7 for the fibre $(V_{x},V_{x}^{+})$:
    $\bigl.\mathcal{L}_{V_{U},V_{U}^{+}}(z)\bigr|_{(x,\chi)}\;=\;c(x,\chi)\cdot\log^{\ast}_{V_{x}(j)}\bigl(z_{x,\chi}\bigr),$
    where $z_{x,\chi}$ is the image of $z$ in $H^{1}(\mathbb{Q}_{p},V_{x}^{\ast}(1)\otimes\chi^{-1})$ and $\log^{\ast}_{V_{x}(j)}$ is the Bloch–Kato dual exponential or logarithm for $V_{x}(j)$.
In particular, after fixing the normalisation of $\eta_{U}$ once and for all, the map $\mathcal{L}_{V_{U},V_{U}^{+}}$ is uniquely determined and interpolates the local Bloch–Kato maps at all classical points of $U$.
Proof.
By Lemma 9, the vector–valued map $\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}$ exists and is AA–linear and $\Lambda(\Gamma)$–linear. Evaluating at the fixed nowhere-vanishing section $\eta_{U}$ gives the scalar–valued map $\mathcal{L}_{V_{U},V_{U}^{+}}$. Since the action of $\mathcal{H}(\Gamma)$ on $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))$ is by $\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$–linear endomorphisms, and evaluation at $\eta_{U}$ is $\mathcal{O}(U)$–linear and $\mathcal{H}(\Gamma)$–equivariant, $\mathcal{L}_{V_{U},V_{U}^{+}}$ is $\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$–linear, proving (i).
For a rigid point $x\in U$, base change along $A\to\kappa(x)$ identifies the fibre of $\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}$ at $x$ with the classical vector–valued Perrin–Riou map $\widetilde{\mathcal{L}}_{V_{x},V_{x}^{+}}$, by Lemma 9(a) together with Theorem 5, Theorem 4 and the comparison between the $(\varphi,\Gamma)$–module and Galois–cohomology constructions in [24, 15]. Evaluating at the specialisation $\eta_{x}$ of $\eta_{U}$ gives the scalar–valued map $\mathcal{L}_{V_{x},V_{x}^{+}}$. By Theorem 7, its specialisations at characters $\chi$ in the Panchishkin range satisfy the stated interpolation property with the Bloch–Kato dual exponentials and logarithms. The additional scalar $c(x,\chi)$ arises from the choice of the basis $\eta_{U}$ of the rank–$1$ line $\mathscr{D}_{\mathrm{cris},U}^{-}$. Since $\eta_{U}$ is fixed once and for all, each $c(x,\chi)$ is non-zero and depends only on this normalisation, as required for (ii). ∎
Remark 4.
The construction above is formally analogous to Hansen’s family–valued regulator on the eigencurve [26, Thm. 1.2.2], where the role of $\mathscr{D}_{\mathrm{cris},U}^{-}$ is played by the line bundle of crystalline periods $D^{\ast}_{\mathrm{cris}}$ and $\widetilde{\mathcal{L}}_{V_{U},V_{U}^{+}}$ is the “Log” map. The only difference is that in our setting $V_{U}$ is a 44–dimensional Rankin–Selberg representation with an rr–dimensional Panchishkin subspace, but the $(\varphi,\Gamma)$–module arguments are identical, thanks to [19, 20].
(3.4) Beilinson–Flach Euler system in the universal deformation family.
We now recall the Euler system input and its extension to the universal deformation setting.
Theorem 10 (Rankin–Selberg Beilinson–Flach Euler system).
Let $f,g$ be classical cuspidal newforms of weights $\geq 2$ and levels prime to $p$ as in Step 1, and let $V(f),V(g)$ be their associated 22–dimensional pp–adic Galois representations. Then there exists an Euler system of Beilinson–Flach classes
$\bigl\{\mathcal{BF}_{m}(f,g)\bigr\}_{m\geq 1}\subset H^{1}\bigl(\mathbb{Q}(\mu_{m}),V(f)^{\ast}(1)\otimes V(g)^{\ast}(1)\bigr)$
satisfying the usual norm–compatibility relations away from $p$, and whose local components at $p$ lie in the Bloch–Kato finite subspaces.