### v2

# Towards a finite–slope universal Rankin–Selberg $p$-adic $L$-function

Abstract
This article studies the finite–slope analogue of Loeffler’s conjectural framework for Rankin–Selberg $p$-adic $L$-functions in universal deformation families. Starting from residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ of tame level $1$ satisfying Hypothesis 3.1 of [13], we consider the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of loc. cit., where the first factor varies in the ordinary Hida deformation and the second factor in the unrestricted universal deformation space.
We fix a classical point $x_{0}$ on a suitable parabolic eigenvariety for which the first factor is non–ordinary but of small finite slope. Using Liu’s global triangulation theorem, the cohomology of families of $(\varphi,\Gamma)$-modules due to Kedlaya–Pottharst–Xiao, and the Perrin–Riou / Loeffler–Zerbes regulator formalism, we attach to the resulting finite–slope Panchishkin data over a neighbourhood $U$ of $x_{0}$ a family “big logarithm”
$\mathcal{L}_{V_{U},V_{U}^{+}}\colon H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)\cong\mathcal{O}(U\times\mathscr{W}),$
which interpolates the Bloch–Kato dual exponentials at all classical points of $U$.
Assuming, in addition, the existence of a global Beilinson–Flach Euler system for this half–ordinary universal deformation family (formulated precisely as Conjecture 2.15 below), we apply this family regulator to the resulting Iwasawa cohomology class and obtain a rigid–analytic function
$L_{p}^{\mathrm{fs}}:=\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})\in\mathcal{O}(U\times\mathscr{W}),$
which we refer to as a finite–slope universal Rankin–Selberg $p$-adic $L$-function. Under standard global and local hypotheses, and conditional on Conjecture 2.15, we show that $L_{p}^{\mathrm{fs}}$ satisfies the expected interpolation formula for all Deligne–critical values of the complex Rankin–Selberg $L$-functions $L(f_{x}\otimes g_{x},s)$ at classical points $(x,\kappa)\in U\times\mathscr{W}$.
Thus the main unconditional output of this work is the construction of a Perrin–Riou regulator for a genuinely finite–slope Panchishkin family over a parabolic eigenvariety in the universal deformation setting. The existence of the associated universal Beilinson–Flach Euler system and the resulting $p$-adic $L$-function remains conjectural; if it holds, our construction would give a finite–slope analogue of the Rankin–Selberg case of Loeffler’s Conjecture 2.8 in [13] for a concrete $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ half–ordinary universal deformation family.
1 Introduction
1.1 Global goal and formulation of the problem
We fix a prime $p\geq 3$ and an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. Let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and let $G_{\mathbb{Q},\{p\}}\subset G_{\mathbb{Q}}$ denote the Galois group of the maximal extension of $\mathbb{Q}$ unramified outside $p$.
Let
$\bar{\rho}_{1},\bar{\rho}_{2}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(\mathbb{F}_{p})$
be continuous, odd, absolutely irreducible Galois representations arising from cuspidal newforms $f_{0},g_{0}$ of weights $k_{1},k_{2}\geq 2$ and tame level $N=1$ (unramified outside $p$), so that Hypothesis 3.1 of [13] applies. In particular, we assume that:
  1. (a)
    both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ are modular of level $1$ and weight at least $2$;
  2. (b)
    the usual Taylor–Wiles hypotheses hold (absolute irreducibility, oddness, and minimality at auxiliary primes);
  3. (c)
    at $p$ the representation $\bar{\rho}_{1}$ is ordinary and admits a fixed ordinary refinement in the sense of [13, Def. 3.11].
We keep this tame level $N=1$ fixed throughout.
Let $R_{1}$ be the ordinary Hida Hecke algebra attached to $\bar{\rho}_{1}$, and let
$\mathscr{X}_{1}:=\operatorname{Spf}(R_{1})^{\operatorname{rig}}$
be the corresponding ordinary eigenvariety, as in [13, §3.2, Prop. 3.14], building on the ordinary families of Hida and the $R=\mathbb{T}$ theorems of Mazur, Böckle and Emerton. Let $R_{2}$ be the universal deformation ring of $\bar{\rho}_{2}$ (for deformations unramified outside $p$) and let $(R_{2},V_{2})$ be the Galois deformation family considered in [13, Def. 3.3, Thm. 3.4] and in [14, 15]. We denote
$\mathscr{X}_{2}:=\operatorname{Spf}(R_{2})^{\operatorname{rig}}$
for the associated rigid analytic space.
Following [13, Ex. 3.17], there is a Panchishkin family $(R,V,V^{+})$ for the Rankin–Selberg tensor
$V:=V_{1}^{\mathrm{ord}}\otimes V_{2}$
over
$R:=R_{1}\widehat{\otimes}_{\mathbb{Z}_{p}}R_{2},$
where $V_{1}^{\mathrm{ord}}$ is the universal ordinary deformation of $\bar{\rho}_{1}$ and $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ is the rank one ordinary submodule at $p$. More precisely, there is a rank $r$ locally free $R$-submodule
$V^{+}\subset V\big|_{G_{\mathbb{Q}_{p}}}$
which is stable under $G_{\mathbb{Q}_{p}}$ and satisfies the rr–Panchishkin condition in the sense of [13, Def. 2.1, Def. 2.3].
[…]
Loeffler’s Conjecture 2.8 in [13] predicts the existence of a rank 0 Euler system (i.e. a $p$-adic $L$-function) attached to $(R,V,V^{+})$. In the ordinary setting this is now known by [15, Thm. 3.5], building on Urban’s construction via nearly overconvergent forms [7] and Hida’s earlier work on Rankin–Selberg $p$-adic $L$-functions [1, 2].
In [13, §5.5] a finite-slope analogue of this conjecture is sketched on the Coleman–Mazur eigencurve, and in [13, §6] related conjectures over big parabolic eigenvarieties are formulated; see also [9, 12]. Apart from very special cases for $\operatorname{GL}_{2}$-families and the rank one case of [11], there is at present no general finite-slope result in the universal deformation setting.
We now formulate the basic question that motivates this work.
Question (finite-slope universal Rankin–Selberg). Let $\mathscr{E}_{1}$ be the Coleman–Mazur eigencurve (or more generally the big parabolic eigenvariety of Barrera Salazar–Williams) attached to $\bar{\rho}_{1}$ [23, 9], and consider the product eigenvariety
$\mathscr{E}:=\mathscr{E}_{1}\times\mathscr{X}_{2}$
together with the natural weight map to the pp–adic weight space $\mathscr{W}$ and the global Galois family $(V,V^{+})$ coming from [13, Ex. 3.17]. Given a classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ of cuspidal eigenforms with $f_{0}$ of non–ordinary finite slope at $p$, does there exist a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on a neighbourhood $U$ of $x_{0}$ whose specialisations interpolate all Deligne–critical Rankin–Selberg values $L(f_{x}\otimes g_{x},s_{x})$ for classical points $x\in U$?
This is a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ instance of the finite-slope variants of [13, Conj. 2.8] over parabolic eigenvarieties. In this paper we give a conditional positive answer, assuming a natural Euler–system conjecture for the relevant universal deformation family.
To treat this question we need a precise description of the eigenvariety $\mathscr{E}$ near $x_{0}$. We recall the relevant geometric input from [9, 29], in a form adapted to our setting.
Lemma 1.1 (Good neighbourhood on the parabolic eigenvariety).
Let $\mathscr{E}$ be the parabolic eigenvariety of [9] with weight space $\mathscr{W}$ and weight map $w:\mathscr{E}\to\mathscr{W}$. Let $x_{0}\in\mathscr{E}$ be a point corresponding to a cuspidal automorphic representation of regular (cohomological) weight, and fix a pp–refinement which is QQ–non-critical in the sense of [9, Def. 5.13]. Assume that the derived group $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series. Then there exists an affinoid neighbourhood $U\subset\mathscr{E}$ of $x_{0}$ such that:
  1. (i)
    $\mathscr{E}$ is smooth over $\mathbb{Q}_{p}$ at every point of $U$;
  2. (ii)
    the restriction $w|_{U}:U\to w(U)\subset\mathscr{W}$ is finite étale;
  3. (iii)
    classical cuspidal points are Zariski dense in $U$.
Proof.
By [9, Def. 5.11] and [9, Rem. 1.4], the overconvergent defect $\ell_{Q}(x_{0})$ vanishes at a QQ–non-critical point; in particular $x_{0}$ lies in the interior locus in the sense of [9, Def. 5.11]. Hence [9, Prop. 5.12] implies that every irreducible component $V$ of $\mathscr{E}$ containing $x_{0}$ has dimension at least $\dim\mathscr{W}$.
On the other hand, $\mathscr{E}$ is constructed as a closed subspace of the universal eigenvariety of Hansen [29]. By [29, Thm. 4.5.1(i)], applied to the corresponding classical, non-critical interior point of the universal eigenvariety, every irreducible component of the latter has dimension equal to the weight space. Since $\mathscr{E}$ is a closed subspace of this universal eigenvariety, every irreducible component $V$ of $\mathscr{E}$ through $x_{0}$ has dimension at most $\dim\mathscr{W}$. Thus $\dim V=\dim\mathscr{W}$ for each component $V$ passing through $x_{0}$.
Under the same hypotheses, [9, Def. 5.13 and Cor. 5.16] show that on each such component $V$ the classical cuspidal points are Zariski dense.
The weight map $w:\mathscr{E}\to\mathscr{W}$ is finite by [9, Thm. 5.4]. Since the source and the target have the same local dimension at $x_{0}$, the morphism $w$ is unramified at $x_{0}$. For finite morphisms between equidimensional rigid analytic spaces, finiteness and unramifiedness imply that $w$ is étale at $x_{0}$ and that $\mathscr{E}$ is smooth at $x_{0}$. The smooth locus of $\mathscr{E}$ and the étale locus of $w$ are open subsets of $\mathscr{E}$, so we may choose an affinoid neighbourhood $U$ of $x_{0}$ contained in their intersection. By the density of classical points on each component through $x_{0}$, the intersection of $U$ with the classical cuspidal locus is Zariski dense in $U$. ∎
Remark 1.2 (On the discrete-series hypothesis in Lemma 1.1).
In our Rankin–Selberg situation we work with
$G=\mathrm{GL}_{2}\times\mathrm{GL}_{2},$
so that
$G_{\mathrm{der}}(\mathbb{R})=\mathrm{SL}_{2}(\mathbb{R})\times\mathrm{SL}_{2}(\mathbb{R}),$
which is well-known to admit discrete series. Thus the hypothesis that $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series in Lemma 1.1 is automatically satisfied in our setting. The only role of this hypothesis is to allow us to invoke the eigenvariety results of [9, 29], in particular the dimension and density statements of [9, Prop. 5.12, Cor. 5.16] and [29, Thm. 4.5.1(i)], in order to construct an affinoid neighbourhood $U$ of $x_{0}$ which is smooth, finite étale over weight space, and has Zariski-dense classical cuspidal locus. All later arguments only use the existence of such a neighbourhood $U$ and do not appeal directly to the discrete-series condition.
1.2 Target main result
We now specialise to the Rankin–Selberg setting of interest. Fix a finite-slope classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ with the following properties:
  1. (a)
    $f_{0}$ is a $p$-stabilised newform of weight $k_{1}\geq 2$ and level prime to $p$, with $U_{p}$–eigenvalue $\alpha_{f_{0}}$ of slope
    $0<v_{p}(\alpha_{f_{0}})<k_{1}-1,$
    so that $f_{0}$ is non-ordinary of small slope at $p$;
  2. (b)
    $g_{0}$ is the specialisation at a classical arithmetic point of the universal deformation family $(R_{2},V_{2})$, of weight $k_{2}\geq 2$, such that $g_{0}$ is ordinary at $p$, or more generally satisfies a suitable Panchishkin condition as in [13, Def. 2.1, Def. 2.3] and [15, §2].
Let $U\subset\mathscr{E}$ be the affinoid neighbourhood of $x_{0}$ given by Lemma 1.1. Thus $\mathscr{E}$ is smooth at every point of $U$, the weight map $w$ is finite étale on $U$, and classical cuspidal points are Zariski dense in $U$.
After shrinking $U$ if necessary, we may and do assume that:
  1. (i)
    the eigenvariety is smooth on $U$ and the weight map $w\colon U\to w(U)\subset\mathscr{W}$ is finite étale (Lemma 1.1);
  2. (ii)
    the slopes of the $U_{p}$-eigenvalues of $f_{x}$ remain $k_{1}(x)-1$ for all $x\in U$, so that overconvergent classicality holds for $f_{x}$ by Coleman [22, Thm. 6.1] (in the form of his Theorem 8.3: weight $k+2$ and slope <k+1<k+1 implies classicality);
  3. (iii)
[…]
The following conjecture is the finite-slope analogue of Loeffler’s universal Rankin–Selberg $p$-adic $L$-function in this neighbourhood.
Conjecture 1.3 (Finite-slope universal Rankin–Selberg $p$-adic $L$-function).
There exists a rigid-analytic function
$L_{p}^{\mathrm{fs}}\in\mathcal{O}(U\times\mathscr{W})$
on $U$ times the $p$-adic cyclotomic weight space $\mathscr{W}$ such that for each classical point $(x,\kappa)\in U\times\mathscr{W}$ corresponding to a pair $(f_{x},g_{x})$ and an integer critical value $s=\kappa(x)$ of $L(f_{x}\otimes g_{x},s)$ in Deligne’s sense, one has an interpolation formula
$L_{p}^{\mathrm{fs}}(x,\kappa)=\frac{E_{p}(f_{x},g_{x},s)}{\Omega_{p}(f_{x},g_{x},\pm)}\cdot\frac{L^{(p)}(f_{x}\otimes g_{x},s)}{(2\pi i)^{2s}},$
where:
  1. (a)
    $E_{p}(f_{x},g_{x},s)$ is the explicit local Euler factor at $p$ appearing in [15, Def. 3.4];
  2. (b)
    $\Omega_{p}(f_{x},g_{x},\pm)$ is a $p$-adic period depending analytically on $x$ (and on the choice of sign $\pm$) and normalised compatibly with [15];
  3. (c)
    $L^{(p)}(f_{x}\otimes g_{x},s)$ is the complex Rankin–Selberg $L$-function with the Euler factor at $p$ omitted.
We now state the main theorem of the paper in a precise special case. It is conditional on an Euler–system conjecture which will be formulated in Conjecture 2.15.
Theorem 1.4 (Main result, conditional on Conjecture 2.15).
Assume:
  1. (H1)
    The residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ satisfy Hypothesis 3.1 of [13] (Taylor–Wiles conditions, local conditions at $p$, and tame level $1$). In particular, both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ arise from cuspidal newforms of level $1$ and weights at least $2$, and we fix once and for all this tame level $N=1$ throughout. Moreover, $\bar{\rho}_{1}$ admits an ordinary refinement at $p$ in the sense of [13, Def. 3.11].
  2. (H2)
    The point $x_{0}$ is crystalline at $p$ and satisfies the small-slope and non-criticality hypotheses needed for overconvergent classicality (in the sense of Coleman [22, Thm. 6.1]).
  3. (H3)
    The Panchishkin condition holds for the local Rankin–Selberg representation at $p$ on $U$ in the sense of [13, Def. 2.1, Def. 2.3] (in particular, the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of loc. cit. uniformly over $U$), so that the triangulation locus of Liu contains $U$.
Assume moreover the Euler–system Conjecture 2.15 for the half–ordinary universal deformation family $(R,V,V^{+})$ of [13, Ex. 3.17]. Then Conjecture 1.3 holds for the neighbourhood $U$. In particular, under (H1)–(H3) and Conjecture 2.15 there exists a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on $U\times\mathscr{W}$ satisfying the above interpolation formula for all classical points in $U$.
The rest of the paper is devoted to the construction of the family regulator $\mathcal{L}_{V_{U},V_{U}^{+}}$ and, under Conjecture 2.15, to the construction and interpolation properties of $L_{p}^{\mathrm{fs}}$. Each step is explicitly referenced to the corresponding statements in [13, 19, 20, 16, 21, 25, 15, 14] so that the argument can be checked in detail.
1.3 Discussion and relation to previous work
In this subsection we give some context for Theorem 1.4, explain how the finite-slope case fits into existing conjectures, and briefly indicate why it is technically more delicate.
Relation to Loeffler’s conjectural framework and to Hao–Loeffler
In his work on $p$-adic $L$-functions in universal deformation families, Loeffler formulates in [13, Conj. 2.8] a general conjecture predicting the existence and interpolation properties of rank 0 Euler systems (equivalently, $p$-adic $L$-functions) attached to 0-Panchishkin families $(R,V,V^{+})$ over suitable “big parabolic eigenvarieties”. A main example is the half–ordinary Rankin–Selberg tensor considered in [13, Ex. 3.17], which is precisely the global deformation setting considered here.
In the ordinary case, this conjectural picture has now been confirmed for the Rankin–Selberg tensor by recent work of Hao–Loeffler [15]. Building on Urban’s nearly overconvergent three–variable Rankin–Selberg $p$-adic $L$-function [7] and the Beilinson–Flach Euler system of Loeffler–Zerbes [21], they construct a universal Rankin–Selberg $p$-adic $L$-function for an ordinary Hida family tensored with a universal deformation family and prove that it interpolates all Deligne–critical Rankin–Selberg values in this setting, together with a functional equation. Their construction is analytic and does not rely on the existence of a universal Euler system.
By contrast, in this work we fix a classical point $x_{0}=(f_{0},g_{0})$ at which the $f_{0}$-factor has non–ordinary finite slope and we work in a neighbourhood $U$ of $x_{0}$ in a parabolic eigenvariety in the sense of Barrera Salazar–Williams [9]. Theorem 1.4 should therefore be viewed as a conditional finite-slope analogue of the Rankin–Selberg case of Loeffler’s conjecture over big parabolic eigenvarieties, in a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ setting.
Comparison with existing finite-slope constructions
Finite-slope $p$-adic $L$-functions attached to families of automorphic forms have been constructed in several other settings, but usually not in the universal deformation framework.
For Rankin–Selberg convolutions of modular forms, Urban’s three–variable $p$-adic $L$-function [7] is constructed on eigenvarieties using nearly overconvergent modular forms and overconvergent cohomology and requires a nearly ordinary hypothesis at $p$. These functions live over eigenvarieties and are not formulated in terms of universal Galois deformation spaces.
The geometric methods of Andreatta–Iovita and their collaborators yield triple product $p$-adic $L$-functions for finite-slope families of modular forms via the theory of overconvergent sheaves and the spectral halo [11]. Here again the base spaces are eigenvarieties and the Galois representations are not packaged as a single Panchishkin family over a universal deformation ring.
Barrera Salazar–Dimitrov–Williams construct finite–slope $p$-adic $L$-functions for Shalika families on $\operatorname{GL}_{2n}$ over parabolic eigenvarieties and relate their existence to the local geometry of these eigenvarieties [12]. Their methods are adapted to the standard representation of $\operatorname{GL}_{2n}$ and do not provide a universal deformation interpretation for Rankin–Selberg tensors of $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$.
In contrast, the present article maintains the universal deformation viewpoint of [13, 15], starting from the half–ordinary Panchishkin family $(R,V,V^{+})$ of [13, Ex. 3.17], and then passing to a finite-slope neighbourhood $U$ on the parabolic eigenvariety through the identification of an ordinary locus in $U$ with an affinoid subdomain of $\operatorname{Spf}(R)^{\operatorname{rig}}$. The function $L_{p}^{\mathrm{fs}}$ constructed under Conjecture 2.15 simultaneously interpolates the critical Rankin–Selberg values for all classical points in $U$ and specialises, at ordinary points, to the ordinary universal Rankin–Selberg $p$-adic $L$-function of [15]. In this way it provides a conceptual bridge between the ordinary universal results and finite-slope eigenvariety constructions in the spirit of Urban and Andreatta–Iovita.
Why the finite-slope case is difficult
From the perspective of [13], it is natural to expect that a finite-slope analogue of Conjecture 2.8 should hold whenever one can attach a $(\varphi,\Gamma)$-module to the local Galois representation in a family and construct an appropriate Perrin–Riou regulator. However, there are several substantial obstacles which have so far prevented a general theory, and which in the Rankin–Selberg case are addressed here.
[…]
Second, even after triangulation, one must control the $(\varphi,\Gamma)$- and Iwasawa cohomology of $D_{U}$ and $D_{U}^{+}$ in families, with good base-change and perfectness properties, in order to set up a Perrin–Riou type regulator over the affinoid algebra $\mathcal{O}(U)$. This is achieved here by systematically using the cohomological machinery of Kedlaya–Pottharst–Xiao [19], which guarantees that the relevant cohomology complexes lie in the perfect derived category and behave well under specialisation.
Third, existing constructions of big logarithms and regulators in families (following Perrin–Riou, Coleman, Loeffler–Zerbes, Nakamura, Pottharst, and others; see for instance [18, 21, 16]) typically treat a single de Rham representation or work under an ordinarity hypothesis. The map
$\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$
constructed in this paper is, to the author’s knowledge, one of the first instances of a Perrin–Riou regulator attached to a genuinely finite-slope Panchishkin family over a higher-dimensional parabolic eigenvariety in the universal deformation setting.
Finally, to obtain a $p$-adic $L$-function one needs an Euler system which specialises correctly along $U$ and lies in the critical Iwasawa cohomology group on which $\mathcal{L}_{V_{U},V_{U}^{+}}$ is defined. This compatibility is crucial in order to define $L_{p}^{\mathrm{fs}}$ as $\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})$ and deduce its interpolation formula from explicit reciprocity laws for Beilinson–Flach elements. At present the existence of such a global class $\mathcal{BF}_{U}$ for the half–ordinary universal deformation family is conjectural (Conjecture 2.15); in the ordinary setting [15] construct the universal Rankin–Selberg $p$-adic $L$-function analytically, without passing through a universal Euler system.
Future directions
The techniques developed here suggest several directions for further work, which we only mention briefly.
One would like to relax the crystalline small-slope and non-criticality hypotheses at $p$. This would require a finer analysis of triangulations and local Galois representations at critical slope, possibly using spectral-halo techniques and overconvergent cohomology as in [11, 9].
It should be possible to adapt the strategy of the later sections to other instances of Loeffler’s conjectural framework, for example to Shalika families on $\operatorname{GL}_{2n}$ or to $\operatorname{GSp}_{4}$-type Rankin–Selberg convolutions, combining parabolic eigenvarieties [9, 12] with existing Euler systems and regulators in those settings.
Assuming the Euler-system conjecture, it would be natural to investigate Iwasawa-theoretic applications and formulate main conjectures relating characteristic ideals of universal Selmer modules over $R$ to the ideal generated by $L_{p}^{\mathrm{fs}}$.
Finally, one expects a functional equation for $L_{p}^{\mathrm{fs}}$ relating its values (or derivatives) at critical points to those of other $p$-adic $L$-functions, and one might hope to extract finer arithmetic invariants such as $p$-adic heights and regulators on motivic cohomology classes varying in families.
2 Proof and results
In this section we explain how the hypotheses (H1)–(H3) lead to the construction of the family regulator and, under Conjecture 2.15, to the $p$-adic $L$-function $L_{p}^{\mathrm{fs}}$ of Theorem 1.4. The discussion is divided into three steps.
2.1 Step 1: Global deformation space and half–ordinary Panchishkin family
We briefly recall the deformation-theoretic input and the construction of the half–ordinary Panchishkin family $(R,V,V^{+})$, following [13, 15]. Recall that $p>2$ is fixed, and that we have two absolutely irreducible residual Galois representations
$\bar{\rho}_{i}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(\mathbb{F}),\qquad i=1,2,$
arising from normalised cuspidal Hecke eigenforms $f_{0}$ and $g_{0}$ of tame level $1$ (or more generally of fixed tame level prime to $p$), with $\bar{\rho}_{1}$ ordinary at $p$ and $\bar{\rho}_{2}$ subject to the usual Taylor–Wiles hypotheses. We write $\mathcal{O}$ for a finite extension of $\mathbb{Z}_{p}$ containing the Hecke eigenvalues of $f_{0}$ and $g_{0}$, and we let $\mathrm{CNL}_{\mathcal{O}}$ denote the category of complete Noetherian local $\mathcal{O}$-algebras with residue field $\mathbb{F}$.
2.1.1 Ordinary universal deformation ring for $\bar{\rho}_{1}$
Let $R^{\mathrm{ord}}(\bar{\rho}_{1})$ denote the universal deformation ring parametrising ordinary deformations of $\bar{\rho}_{1}$ as a $G_{\mathbb{Q},\{p\}}$-representation in the sense of Hida and Mazur: for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation
$\rho_{1,A}:G_{\mathbb{Q},\{p\}}\to\operatorname{GL}_{2}(A)$
is called ordinary if its restriction to $G_{\mathbb{Q}_{p}}$ fits into a short exact sequence
$0\longrightarrow V^{\mathrm{ord},+}_{1,A}\longrightarrow V^{\mathrm{ord}}_{1,A}\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V^{\mathrm{ord},-}_{1,A}\longrightarrow 0$
where $V^{\mathrm{ord},+}_{1,A}$ is a rank one direct summand on which $G_{\mathbb{Q}_{p}}$ acts by an unramified character lifting the $p$-adic unit root of the Hecke polynomial of $f_{0}$.
By Mazur’s deformation theory together with Hida theory and the work of Böckle and Emerton (see for example [13, §3.2, Prop. 3.14 and Thm. 3.4]), this ordinary deformation problem is representable by a complete Noetherian local $\mathcal{O}$-algebra $R^{\mathrm{ord}}(\bar{\rho}_{1})$, and there exists a universal ordinary deformation
$\rho_{1}^{\mathrm{ord}}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}\big(R^{\mathrm{ord}}(\bar{\rho}_{1})\big)$
with underlying free rank two $R^{\mathrm{ord}}(\bar{\rho}_{1})$-module $V_{1}^{\mathrm{ord}}$. We denote by $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ the rank one direct summand giving the ordinary filtration at $p$, and by $V_{1}^{\mathrm{ord},-}$ the corresponding quotient.
The ring $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is finite flat over the weight space and is canonically isomorphic to the localised ordinary Hecke algebra $T^{\mathrm{ord}}(\bar{\rho}_{1})$ acting on ordinary $p$-adic modular forms of tame level $1$; in particular $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is reduced and equidimensional of relative dimension $1$ over $\mathcal{O}$ (see [13, §3.2, Prop. 3.14]). For brevity we set
$R_{1}:=R^{\mathrm{ord}}(\bar{\rho}_{1}).$
2.1.2 The unrestricted universal deformation ring and representation for $\bar{\rho}_{2}$
For the second factor, we do not impose any ordinary or nearly ordinary local condition at $p$. We let $R(\bar{\rho}_{2})$ be the universal deformation ring parametrising deformations of $\bar{\rho}_{2}$ as a $G_{\mathbb{Q},\{p\}}$-representation (unramified outside $p$), as in [13, Def. 3.3]. Thus, for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation to $A$ is simply a continuous representation
$G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(A)$
lifting $\bar{\rho}_{2}$ and unramified outside $p$.
The main $R=\mathbb{T}$ theorem of Böckle–Emerton in this context (as formulated in [13, Thm. 3.4]) shows that:
  1. (a)
    The ring $R(\bar{\rho}_{2})$ is a reduced complete intersection ring, flat over $\mathcal{O}$ of relative dimension $3$.
  2. (b)
    There is a canonical isomorphism
    $R(\bar{\rho}_{2})\cong T(\bar{\rho}_{2}),$
    where $T(\bar{\rho}_{2})$ is the localisation at the maximal ideal corresponding to $\bar{\rho}_{2}$ of the prime-to-$p$ Hecke algebra acting on the space $S(1,\mathcal{O})$ of cuspidal $p$-adic modular forms of tame level $1$.
By definition of the universal deformation, there is a free rank two $R(\bar{\rho}_{2})$-module $V_{2}$ equipped with a continuous $G_{\mathbb{Q},\{p\}}$-action lifting $\bar{\rho}_{2}$; we regard this as the universal $p$-adic Galois representation of type $\bar{\rho}_{2}$. We henceforth abbreviate
$R_{2}:=R(\bar{\rho}_{2}).$
2.1.3 The half–ordinary Rankin–Selberg Panchishkin family
We now combine the two deformation problems. Consider the completed tensor product
$R:=R_{1}\widehat{\otimes}_{\mathcal{O}}R_{2}.$
Since $R_{1}$ and $R_{2}$ are flat complete Noetherian local $\mathcal{O}$-algebras of relative dimensions $1$ and $3$ respectively, their completed tensor product $R$ is again a flat complete Noetherian local $\mathcal{O}$-algebra of relative dimension
$\dim R=\dim R_{1}+\dim R_{2}=1+3=4,$
and the residual representation is $\bar{\rho}_{1}\otimes\bar{\rho}_{2}$.
On $R$ we have the rank four $R$-module
$V:=V_{1}^{\mathrm{ord}}\widehat{\otimes}_{\mathcal{O}}V_{2}$
with its diagonal $G_{\mathbb{Q}}$-action, and the rank two $R$-submodule
$V^{+}:=V_{1}^{\mathrm{ord},+}\widehat{\otimes}_{\mathcal{O}}V_{2}\subset V\big|_{G_{\mathbb{Q}_{p}}},$
which is stable under $G_{\mathbb{Q}_{p}}$ because both factors are. This gives a $G_{\mathbb{Q}_{p}}$-stable filtration
$0\longrightarrow V^{+}\longrightarrow V\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V^{-}\longrightarrow 0$
with $V^{+}$ and $V^{-}$ of rank $2$ over $R$.
This is exactly the datum considered in [13, Ex. 3.17], specialised to the case where the first factor is already ordinary. In the notation of loc. cit., we are taking
$R=R^{\mathrm{ord}}(\bar{\rho}_{1})\widehat{\otimes}_{\mathcal{O}}R(\bar{\rho}_{2}),\qquad V=V_{1}^{\mathrm{ord}}\widehat{\otimes}_{\mathcal{O}}V_{2},\qquad V^{+}=V_{1}^{\mathrm{ord},+}\widehat{\otimes}_{\mathcal{O}}V_{2}.$
Loeffler proves in [13, Ex. 3.17] that $(V,V^{+})$ is a 0-Panchishkin family over $R$ in his sense: for every arithmetic point $\kappa$ of $\operatorname{Spf}(R)^{\operatorname{rig}}$ corresponding to a pair $(f,g)$ of classical eigenforms, the specialisation $(V_{\kappa},V_{\kappa}^{+})$ coincides with the Rankin–Selberg Galois representation $V(f)\otimes V(g)$ together with the usual half–ordinary subspace at $p$. Moreover, the interpolation set
[…]
is Zariski-dense in $\operatorname{Spf}(R)^{\operatorname{rig}}$. Thus the triple $(R,V,V^{+})$ provides the global deformation space and Panchishkin data referred to in Theorem 1.4.
2.1.4 Relation with the eigenvariety and with the ordinary locus in $U$
We now explain how the universal half–ordinary Panchishkin family of [13, Ex. 3.17] interacts with the parabolic eigenvariety $\mathscr{E}$ and with the neighbourhood $U$ of our fixed point $x_{0}$.
Let $E$ denote the parabolic eigenvariety of [9] attached to $G=\operatorname{GL}_{2}\times\operatorname{GL}_{2}$, with weight space $\mathcal{W}$ and weight map $w:E\to\mathcal{W}$. We keep the notation
$X_{1}:=\operatorname{Spf}(R_{1})^{\operatorname{rig}},\qquad X_{2}:=\operatorname{Spf}(R_{2})^{\operatorname{rig}},\qquad X:=X_{1}\times X_{2}=\operatorname{Spf}(R)^{\operatorname{rig}}.$
By construction, $X_{1}$ dominates the ordinary locus of the Coleman–Mazur eigencurve (and more generally the parabolic eigenvariety) attached to $\bar{\rho}_{1}$; this is a consequence of Hida theory together with the $R=\mathbb{T}$ theorems used in [13, §3.2]. The space $X_{2}$ is the universal deformation space for $\bar{\rho}_{2}$.
On the locus where the first factor is ordinary at $p$ there is a natural morphism
$X=X_{1}\times X_{2}\longrightarrow E$
whose points correspond to pairs of classical cusp forms $(f,g)$ with $f$ ordinary at $p$, together with an ordinary refinement on the $f$-factor. On this locus the specialisation $(V_{x},V_{x}^{+})$ of $(V,V^{+})$ is the Rankin–Selberg representation $V(f)\otimes V(g)$ together with the usual half–ordinary subspace at $p$, as in [13, Ex. 3.17] and [15].
Let $E^{\mathrm{ord}}\subset E$ denote the ordinary locus for the first factor, and set
$U^{\mathrm{ord}}:=U\cap E^{\mathrm{ord}}.$
This is an admissible open subset of $U$. Since ordinary classical points are Zariski dense in the ordinary eigenvariety and $U$ contains a Zariski-dense set of classical points, we may, after shrinking $U$ slightly if necessary, assume that $U^{\mathrm{ord}}$ is non-empty and that classical points are still Zariski dense in $U^{\mathrm{ord}}$.
On $E^{\mathrm{ord}}$ the eigenvariety machine of [9, 29], together with the Rankin–Selberg construction of [15], shows that the map $X\to E$ above identifies an affinoid neighbourhood of any classical half–ordinary point with an affinoid subdomain of $X$ (see [9, Thm. 5.4], [13, §3.2] and [15, §2]). In particular, there exists an affinoid subdomain
$U_{R}\subset X$
and an isomorphism of rigid spaces over weight space
$\iota:U^{\mathrm{ord}}\xrightarrow{\ \sim\ }U_{R}\subset X$
such that, for every classical point $x\in U^{\mathrm{ord}}$ corresponding to a pair $(f_{x},g_{x})$ with $f_{x}$ ordinary at $p$, the specialisation of the universal family $(V,V^{+})$ at $\iota(x)$ coincides with the Rankin–Selberg representation
$V(f_{x})\otimes V(g_{x})$
together with its standard half–ordinary subspace at $p$.
Two points are important here.
First, the finite-slope point $x_{0}=(f_{0},g_{0})$ does not lie in the image of $X\to E$ when $f_{0}$ is non-ordinary at $p$. Thus $x_{0}$ itself does not belong to $U^{\mathrm{ord}}$, and we do not try to realise $x_{0}$ as a point of the rigid fibre of $R$.
[…]
2.2 Step 2: Global triangulation and perfect cohomology complexes
In this step we work entirely over the fixed affinoid neighbourhood $U\subset E$ of $x_{0}$. We recall the family of local Galois representations on $U$, attach to it a family of $(\varphi,\Gamma)$-modules, and record the finiteness and base-change properties of the associated cohomology complexes, following [20, 19].
2.2.1 The relative $(\varphi,\Gamma)$-module
By the eigenvariety machine of Hansen and Barrera Salazar–Williams (see [29, Thm. 4.5.1] and [9, §5]), there exists a locally free rank four $\mathcal{O}(U)$-module $V_{U}$ equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$, unramified outside $p$, such that for every classical point $x\in U$ corresponding to a pair of cuspidal eigenforms $(f_{x},g_{x})$ the fibre $V_{x}$ is canonically isomorphic to the Rankin–Selberg tensor
$V(f_{x})\otimes V(g_{x}).$
The $U_{p}$-eigenvalues on $U$ give an analytic function $\alpha:U\to\mathcal{O}(U)^{\times}$ encoding a refinement of $V_{U}\big|_{G_{\mathbb{Q}_{p}}}$; this makes $V_{U}\big|_{G_{\mathbb{Q}_{p}}}$ into a weakly refined family in the sense of [20, Def. 1.5], cf. the discussion in [20, §5] and [29, §4.5].
Let
$\mathscr{D}:=D_{\operatorname{rig}}^{\dagger}\bigl(V_{U}\big|_{G_{\mathbb{Q}_{p}}}\bigr)$
be the associated family of $(\varphi,\Gamma)$-modules over the relative Robba ring $\mathcal{R}_{\mathcal{O}(U)}$; cf. [19, Thm. 2.2.17]. For any point $x\in U$ we write $\mathscr{D}_{x}$ for the fibre; by construction
$\mathscr{D}_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x})$
as a $(\varphi,\Gamma)$-module over the usual Robba ring over $\mathbb{Q}_{p}$.
Remark 2.1.
The existence, functoriality and base-change properties of $D_{\operatorname{rig}}^{\dagger}$ for families of $p$-adic Galois representations are given by [19, Thm. 2.2.17]. Concretely, if $A$ is a $\mathbb{Q}_{p}$-affinoid algebra and $V_{A}$ is a finite projective $A$-module with a continuous $A$-linear action of $G_{\mathbb{Q}_{p}}$, then there is an associated $(\varphi,\Gamma)$-module $D_{\operatorname{rig}}^{\dagger}(V_{A})$ over the relative Robba ring $\mathcal{R}_{A}$, and for any morphism $A\to B$ of $\mathbb{Q}_{p}$-affinoid algebras one has a natural base-change isomorphism
$D_{\operatorname{rig}}^{\dagger}(V_{A})\widehat{\otimes}_{A}B\cong D_{\operatorname{rig}}^{\dagger}(V_{A}\otimes_{A}B).$
2.2.2 Global triangulation à la Liu
We recall Liu’s global triangulation theorem in the form needed here.
Theorem 2.2 (Liu, global triangulation).
Let $X$ be a reduced separated rigid analytic space over $\mathbb{Q}_{p}$, and let $V_{X}$ be a weakly refined family of $p$-adic representations of $G_{\mathbb{Q}_{p}}$ over $X$ in the sense of [20, Def. 1.5]. Let $\mathscr{D}_{X}:=D_{\operatorname{rig}}^{\dagger}(V_{X})$ be the associated family of $(\varphi,\Gamma)$-modules. Then there exists a Zariski open and dense subspace $X^{\mathrm{tri}}\subset X$ such that $\mathscr{D}_{X}\big|_{X^{\mathrm{tri}}}$ admits a global triangulation: that is, a filtration by $(\varphi,\Gamma)$-submodules
$0=\operatorname{Fil}^{0}\mathscr{D}_{X}\subset\operatorname{Fil}^{1}\mathscr{D}_{X}\subset\cdots\subset\operatorname{Fil}^{d}\mathscr{D}_{X}=\mathscr{D}_{X}$
whose graded pieces $\operatorname{gr}^{i}\mathscr{D}_{X}:=\operatorname{Fil}^{i}/\operatorname{Fil}^{i-1}$ are rank one $(\varphi,\Gamma)$-modules with prescribed parameters. Moreover, the locus of global triangulation contains all regular non-critical points of $X$ [20, Thm. 1.8].
We apply Theorem 2.2 with $X=U$ and $V_{X}=V_{U}\big|_{G_{\mathbb{Q}_{p}}}$. By hypotheses (H2) and (H3), the point $x_{0}$ is crystalline with small slope and the local conditions at $p$ define a non-critical Panchishkin situation in the sense of [20, Def. 5.29]; in particular $x_{0}$ is a regular non-critical point. Hence there is a Zariski open neighbourhood $U^{\mathrm{tri}}\subset U$ of $x_{0}$ on which $\mathscr{D}$ admits a global triangulation.
After shrinking $U$ if necessary, we may and do assume that $U\subset U^{\mathrm{tri}}$ and that there exists a filtration
$0=\operatorname{Fil}^{0}\mathscr{D}\subset\operatorname{Fil}^{1}\mathscr{D}\subset\operatorname{Fil}^{2}\mathscr{D}\subset\operatorname{Fil}^{3}\mathscr{D}\subset\operatorname{Fil}^{4}\mathscr{D}=\mathscr{D}$
of $\mathscr{D}$ by $(\varphi,\Gamma)$-submodules such that each $\operatorname{gr}^{i}\mathscr{D}$ is locally free of rank one over the relative Robba ring over $U$.
Definition 2.3.
Over the neighbourhood $U$ we define
$\mathscr{D}^{+}:=\operatorname{Fil}^{r}\mathscr{D},$
[…]
Proposition 2.4 (Compatibility with the Panchishkin local condition).
[…]
$(\mathscr{D}^{+})_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x}^{\mathrm{ord},+}),$
where $V_{x}^{\mathrm{ord},+}\subset V_{x}$ is the rank $r$ ordinary $G_{\mathbb{Q}_{p}}$-subrepresentation coming from the Hida deformation of $\bar{\rho}_{1}$.
Proof.
[…]
2.2.3 Finiteness and base change for $(\varphi,\Gamma)$-cohomology
We now record the finiteness and base-change properties of the $(\varphi,\Gamma)$-cohomology and Iwasawa cohomology of $\mathscr{D}$ and $\mathscr{D}^{+}$ over $U$, following [19]. Let us briefly recall the complexes used by Kedlaya–Pottharst–Xiao. For a $(\varphi,\Gamma)$-module $\mathsf{M}$ over a relative Robba ring $\mathcal{R}_{A}$, we denote by $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ the complex computing the usual $(\varphi,\Gamma)$-cohomology, and by $C^{\bullet}_{\psi}(\mathsf{M})$ the complex computing Iwasawa cohomology, as defined in [19, §4].
Theorem 2.5 (Kedlaya–Pottharst–Xiao).
Let $A$ be an affinoid $\mathbb{Q}_{p}$-algebra and let $\mathsf{M}$ be a $(\varphi,\Gamma)$-module over the relative Robba ring $\mathcal{R}_{A}$. Then:
  1. (i)
    The Iwasawa cohomology complex $C^{\bullet}_{\psi}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{A}^{\infty}(\Gamma)\bigr)$, and the $(\varphi,\Gamma)$-cohomology complex $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}(A)$ [19, Thm. 4.4.1, Thm. 4.4.2].
  2. (ii)
    For any morphism of affinoid algebras $A\to B$, the natural maps
    $C^{\bullet}_{\psi}(\mathsf{M})\otimes^{\mathbf{L}}_{\mathcal{R}_{A}^{\infty}(\Gamma)}\mathcal{R}_{B}^{\infty}(\Gamma)\longrightarrow C^{\bullet}_{\psi}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr),$
    $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})\otimes^{\mathbf{L}}_{A}B\longrightarrow C^{\bullet}_{\varphi,\Gamma}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr)$
    are quasi-isomorphisms [19, Thm. 4.4.3].
We apply Theorem 2.5 with $A=\mathcal{O}(U)$ and $\mathsf{M}=\mathscr{D}$ or $\mathsf{M}=\mathscr{D}^{+}$. This shows that
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}),\ C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}^{+})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{O}(U)\bigr),$
while
$C^{\bullet}_{\psi}(\mathscr{D}),\ C^{\bullet}_{\psi}(\mathscr{D}^{+})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{\mathcal{O}(U)}^{\infty}(\Gamma)\bigr).$
Moreover, for any morphism of affinoid algebras $\mathcal{O}(U)\to B$ (for instance $B=\mathcal{O}(U^{\prime})$ for an affinoid subdomain $U^{\prime}\subset U$, or $B=k(x)$ for the residue field at a point $x\in U$), the base-change isomorphisms of Theorem 2.5(ii) give quasi-isomorphisms
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{O}(U)}B\xrightarrow{\ \sim\ }C^{\bullet}_{\varphi,\Gamma}\bigl(\mathscr{D}\widehat{\otimes}_{\mathcal{O}(U)}B\bigr),$
and similarly for $\mathscr{D}^{+}$, as well as
$C^{\bullet}_{\psi}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{R}_{\mathcal{O}(U)}^{\infty}(\Gamma)}\mathcal{R}_{B}^{\infty}(\Gamma)\xrightarrow{\ \sim\ }C^{\bullet}_{\psi}\bigl(\mathscr{D}\widehat{\otimes}_{\mathcal{O}(U)}B\bigr),$
and analogously for $\mathscr{D}^{+}$. In particular, the formation of these complexes is compatible with restriction to smaller affinoid neighbourhoods and with specialisation at points of $U$.
Remark 2.6.
[…]
$\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Z}_{p}}\mathcal{O}(U),$
and to relate its specialisations at classical points to the Bloch–Kato exponentials.
2.3 Step 3: Construction of the big logarithm and the finite-slope regulator
In this final step we fix once and for all the affinoid neighbourhood $U$ and the triangulated $(\varphi,\Gamma)$-module
$(\mathscr{D},\mathscr{D}^{+})$
over $U$ constructed in Step 2. Thus $V_{U}$ is the rank four finite projective $\mathcal{O}(U)$-module equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$ whose specialisations are the Rankin–Selberg Galois representations $V_{x}$, and
$\mathscr{D}=D_{\operatorname{rig}}^{\dagger}\bigl(V_{U}\big|_{G_{\mathbb{Q}_{p}}}\bigr)$
[…]
We first introduce Iwasawa-theoretic notation and recall the local Perrin–Riou big logarithm for a single Panchishkin representation. We then explain how to extend it to the family $(V_{U},V_{U}^{+})$ using relative $(\varphi,\Gamma)$-modules and the Perrin–Riou formalism, and finally we state the Euler–system conjecture and the resulting conditional construction of $L_{p}^{\mathrm{fs}}$.
2.3.1 Iwasawa cohomology and the distribution algebra
Let
$\Gamma:=\operatorname{Gal}\bigl(\mathbb{Q}_{p}(\mu_{p^{\infty}})/\mathbb{Q}_{p}\bigr)\cong\operatorname{Gal}\bigl(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}\bigr),\qquad\Lambda(\Gamma):=\mathbb{Z}_{p}[[\Gamma]].$
Let
$\mathcal{H}(\Gamma):=\mathcal{H}_{\mathbb{Q}_{p}}(\Gamma)$
denote the algebra of $\mathbb{Q}_{p}$-valued locally analytic distributions on $\Gamma$, i.e. the completion of $\Lambda_{\mathbb{Q}_{p}}(\Gamma)$ in its natural Fréchet topology (cf. [30, §2.2] and [31, §2.1]). Equivalently, $\mathcal{H}(\Gamma)$ is the strong continuous dual of the space of locally analytic $\mathbb{Q}_{p}$-valued functions on $\Gamma$.
For any $p$-adic representation $W$ of $G_{\mathbb{Q}_{p}}$ on a finite-dimensional $\mathbb{Q}_{p}$-vector space, its (cyclotomic) Iwasawa cohomology is defined by
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W):=H^{i}\bigl(\mathbb{Q}_{p},W\otimes_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
where $\Lambda(\Gamma)^{\vee}:=\operatorname{Hom}_{\mathrm{cts}}(\Lambda(\Gamma),\mathbb{Q}_{p})$ with the natural $G_{\mathbb{Q}_{p}}$-action; see [18, App. A.2–A.4] and Greenberg [28]. By construction, $\Lambda(\Gamma)$ acts on $\Lambda(\Gamma)^{\vee}$ via the right regular representation, and hence $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ carries a natural $\Lambda(\Gamma)$-module structure for every $i$.
Lemma 2.7 (Finiteness of local Iwasawa cohomology).
If $W$ is a finite-dimensional $\mathbb{Q}_{p}$-representation of $G_{\mathbb{Q}_{p}}$, then $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ is a finitely generated $\Lambda(\Gamma)$-module for all $i\geq 0$. In particular, $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ is finitely generated over $\Lambda(\Gamma)$.
Proof.
Choose a $G_{\mathbb{Q}_{p}}$-stable $\mathbb{Z}_{p}$-lattice $T\subset W$, and write $K_{\infty}=\mathbb{Q}_{p}(\mu_{p^{\infty}})$. Following [18, App. A.2], let
[…]
with respect to the corestriction maps; then $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is naturally a $\Lambda(\Gamma)$-module, and there is a canonical identification
$Z^{i}_{\infty}(\mathbb{Q}_{p},T)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},T):=H^{i}\bigl(\mathbb{Q}_{p},T\otimes_{\mathbb{Z}_{p}}\Lambda(\Gamma)^{\vee}\bigr)$
for all $i$ (see [18, App. A.2–A.3]). Moreover, Proposition A.2.3 of loc. cit. shows that $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is a finitely generated $\Lambda(\Gamma)$-module for $i=0,1,2$, and $Z^{i}_{\infty}(\mathbb{Q}_{p},T)=0$ for $i\geq 3$. Since $W=T\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$, we have
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\cong Z^{i}_{\infty}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$
as $\Lambda(\Gamma)$-modules, and the finite generation of $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ implies the finite generation of $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ for all $i$. ∎
In order to apply Perrin–Riou’s regulator, we pass from $\Lambda(\Gamma)$-coefficients to the algebra $\mathcal{H}(\Gamma)$ of locally analytic distributions by scalar extension along the canonical map $\Lambda(\Gamma)\to\mathcal{H}(\Gamma)$ and set
$H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)_{\mathcal{H}}:=H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{H}(\Gamma).$
When no confusion can arise, we continue to denote this $\mathcal{H}(\Gamma)$-module simply by $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$.
If $A$ is a $\mathbb{Q}_{p}$-affinoid algebra and $W_{A}$ is a finite projective $A$-module with a continuous $A$-linear $G_{\mathbb{Q}_{p}}$-action, we define the family Iwasawa cohomology by
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A}):=H^{i}\bigl(\mathbb{Q}_{p},W_{A}\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
which is a priori an $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$-module. Let $\mathcal{R}_{A}^{\infty}(\Gamma)$ denote the Fréchet–Stein Iwasawa algebra of $\Gamma$ over $A$ considered in [19, §4.2]; there is a natural finite flat homomorphism
$A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)\longrightarrow\mathcal{R}_{A}^{\infty}(\Gamma).$
The relative $(\varphi,\Gamma)$-module formalism of Kedlaya–Pottharst–Xiao gives a more precise description.
Proposition 2.8 (Relative local Iwasawa cohomology via $(\varphi,\Gamma)$-modules).
Let $A$ be a reduced $\mathbb{Q}_{p}$-affinoid algebra and $W_{A}$ a finite projective $A$-module with a continuous $A$-linear $G_{\mathbb{Q}_{p}}$-action. Let $\mathscr{D}_{A}:=D_{\operatorname{rig}}^{\dagger}(W_{A})$ be the associated family of $(\varphi,\Gamma)$-modules over the relative Robba ring $\mathcal{R}_{A}$. Then:
  1. (a)
    there exist functorial complexes
    $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})\in D^{-}_{\mathrm{perf}}(A),\qquad C^{\bullet}_{\psi}(\mathscr{D}_{A})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{A}^{\infty}(\Gamma)\bigr)$
    such that their cohomology groups compute the Galois and Iwasawa cohomology of $W_{A}$ in the sense that
    $H^{i}\bigl(C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})\bigr)\cong H^{i}(\mathbb{Q}_{p},W_{A}),$
    $H^{i}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{R}_{A}^{\infty}(\Gamma)$
    for all $i\geq 0$;
  2. (b)
    the formation of $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, together with the above isomorphisms, commutes with flat base change in $A$;
  3. (c)
    in particular, $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})$ is a finite projective $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$-module, and for every morphism of affinoids $A\to B$ the canonical map
    $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})\widehat{\otimes}_{A\widehat{\otimes}\Lambda(\Gamma)}B\widehat{\otimes}\Lambda(\Gamma)\xrightarrow{\ \sim\ }H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{B})$
    is an isomorphism.
Proof.
The existence, perfectness, and base-change properties of the complexes $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, as well as the identifications of their cohomology with Galois and Iwasawa cohomology after extension of scalars to $\mathcal{R}_{A}^{\infty}(\Gamma)$, are proved in [19, Rem. 4.3.3, Prop. 4.3.6, Cor. 4.3.7, Prop. 4.3.8, Thm. 4.4.3]. The finite projectivity of $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})$ over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$ and its base-change property then follow from the fact that $\mathcal{R}_{A}^{\infty}(\Gamma)$ is a Fréchet–Stein algebra, finite flat over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$, and that $H^{1}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)$ is a coadmissible (hence finite projective) $\mathcal{R}_{A}^{\infty}(\Gamma)$-module; see [19, Lem. 4.3.4]. ∎
We apply this proposition with $A=\mathcal{O}(U)$ and $W_{A}=V_{U}^{\ast}(1)$, and write
$H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))$
for the resulting family of local Iwasawa cohomology groups, viewed as a finite projective module over $\mathcal{O}(U)\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$.
2.3.2 Classical Perrin–Riou big logarithm
We briefly recall the local Perrin–Riou regulator for a single de Rham Panchishkin representation.
Theorem 2.9 (Perrin–Riou big logarithm).
Let $V$ be a finite-dimensional $\mathbb{Q}_{p}$-vector space with a continuous de Rham $G_{\mathbb{Q}_{p}}$-action, and let $V^{+}\subset V$ be a Panchishkin subspace in the usual sense (cf. Panchishkin [3] or [16]). Then there exists a canonical $\Lambda(\Gamma)$-linear map
$\mathcal{L}_{V,V^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\otimes_{\mathbb{Q}_{p}}D_{\mathrm{cris}}(V)$
with the following interpolation property: for every integer twist $V(j)$ in the Bloch–Kato range determined by $V^{+}$ and every continuous character $\chi:\Gamma\to\overline{\mathbb{Q}}_{p}^{\times}$ of finite order, the specialisation of $\mathcal{L}_{V,V^{+}}$ at $\chi$ is, up to an explicit non-zero scalar depending only on $V$ and $\chi$, the Bloch–Kato dual exponential or logarithm map for $V(j)$.
More precisely, for such $\chi$ there is a commutative diagram
[…]
where $\log^{\ast}_{V(j)}$ is the Bloch–Kato dual exponential or logarithm (depending on the Hodge–Tate weights), and $\operatorname{Fil}^{0}$ is the Hodge filtration on $D_{\mathrm{dR}}(V(j))$.